### v1

# Conditional Independence Test Based on Transport Maps

Abstract
Testing conditional independence between two random vectors given a third is a fundamental and challenging problem in statistics, particularly in multivariate nonparametric settings due to the complexity of conditional structures. We propose a novel framework for testing conditional independence using transport maps. At the population level, we show that two well-defined transport maps can transform the conditional independence test into an unconditional independence test, this substantially simplifies the problem. These transport maps are estimated from data using conditional continuous normalizing flow models. Within this framework, we derive a test statistic and prove its consistency under both the null and alternative hypotheses. A permutation-based procedure is employed to evaluate the significance of the test. We validate the proposed method through extensive simulations and real-data analysis. Our numerical studies demonstrate the practical effectiveness of the proposed method for conditional independence testing.
Key words: Continuous normalizing flow, Deep neural networks, Distance correlation, Independence, Permutation.
1 Introduction
Let $\mathbf{X}\in\mathbb{R}^{d_{\mathbf{x}}}$, $\mathbf{Y}\in\mathbb{R}^{d_{\mathbf{y}}}$, and $\mathbf{Z}\in\mathbb{R}^{d_{\mathbf{z}}}$ be three random vectors. We are interested in testing whether $\mathbf{X}$ and $\mathbf{Y}$ are conditionally independent given $\mathbf{Z}$ (Dawid, 1979), formally defined as:
$H_{0}:\ \mathbf{X}\perp\!\!\!\perp\mathbf{Y}\mid\mathbf{Z},\text{ versus }H_{1}:\mathbf{X}\not\perp\!\!\!\perp\mathbf{Y}\mid\mathbf{Z}.$ (1.1)
Conditional independence is a fundamental concept in statistics and machine learning, playing a crucial role in various fields such as causal inference (Spirtes et al., 1993; Pearl, 2009; Spirtes, 2010; Constantinou and Dawid, 2017) and graphical models (Geiger and Pearl, 1993; Koldanov et al., 2017; Fan et al., 2020). Understanding the conditional independence between random vectors allows researchers to simplify complex models, reduce dimensionality, and enhance interpretability.
Numerous methods have been proposed for testing conditional independence. Parametric approaches impose explicit structural assumptions on the conditional distributions $\mathbf{X}\mid\mathbf{Z}$ and $\mathbf{Y}\mid\mathbf{Z}$. For example, Song (2009) adopted a single-index model paired with the Rosenblatt transformation (Rosenblatt, 1952), while Fan et al. (2020) assumed linearity in the conditional distributions. Though such methods achieve root-$n$ consistency under their parametric constraints, their reliance on strict model specifications renders them vulnerable to misspecification bias, limiting their applicability to complex, non-linear dependencies. Beyond parametric constraints, kernel-based methods employ kernel functions to measure dependence without parametric assumptions. For instance, Zhang et al. (2011) and Huang et al. (2022) derived tests using cross-covariance operator in reproducing kernel Hilbert spaces (RKHS). Wang et al. (2015) introduced conditional distance correlation, where they utilized kernel functions to estimate the empirical form. Cai et al. (2022) formulated conditional independence as mutual independence across transformed variables, estimating dependencies via kernel-based empirical measures. Though these methods avoid strict parametric assumptions, they either incur substantial computational costs from kernel matrix operations or exhibit reduced statistical power against alternatives as the dimension of $\mathbf{Z}$ grows, even at moderate dimensionality (e.g., dimension of $\mathbf{Z}$ larger than 3). Recent methodological advances have introduced modern machine learning techniques for conditional independence testing. For example, Sen et al. (2017) formulated conditional independence test as a classification problem, employing gradient-boosted trees and deep neural networks to perform the classification. Bellot and van der Schaar (2019) and Shi et al. (2021) developed frameworks using generative adversarial networks (GANs), while Duong and Nguyen (2024) proposed a normalizing flow-based method. However, compared to state-of-the-art generative models such as diffusion models and continuous normalizing flows, GANs and normalizing flows often suffer from training instability. Yang et al. (2024) explored the use of diffusion models for testing but restricted their analysis to univariate variable pairs.
In this work, we propose a framework for conditional independence testing based on transport maps. At the population level, we demonstrate that two invertible maps from the conditional distributions to simple reference distributions allow us to formulate conditional independence test as unconditional independence test, thereby simplifying the formulation. We employ conditional continuous normalizing flows and independence measures to effectively address conditional independence testing. Our contributions are summarized as follows:
  1. 1.
    A novel formulation of conditional independence. We demonstrate that testing conditional independence between $\mathbf{X}$ and $\mathbf{Y}$ given $\mathbf{Z}$ can be simplified to testing unconditional independence through invertible transformations. By transforming $(\mathbf{X},\mathbf{Z})$ and $(\mathbf{Y},\mathbf{Z})$ into new variables (e.g., Gaussian variables) independent of $\mathbf{Z}$, conditional independence becomes equivalent to unconditional independence. We rigorously prove this equivalence in Lemma 1. This formulation significantly reduces the complexity associated with conditional independence structure.
  2. 2.
    A model-free test statistic. We introduce a model-free test statistic for conditional independence testing. By using the continuous normalizing flow as a transport map, we can apply popular independence measures, such as distance correlation, which do not depend on specific model assumptions about the data distribution. The consistency guarantees are established in Theorem 5.
  3. 3.
    Performance across dimensions. Through extensive simulations, we demonstrate that our proposed test exhibits superior performance across univariate, multivariate, and moderately high-dimensional pairs $\mathbf{X}$ and $\mathbf{Y}$ (with dimensions $d_{\mathbf{x}}$ and $d_{\mathbf{y}}$ ranging from 1 to 50) in the presence of low or moderately high-dimensional confounders $\mathbf{Z}$ (with dimensions $d_{\mathbf{z}}$ ranging from 2 to 100).
  4. 4.
    Robustness to independence measures. After transforming a conditional independence problem into an unconditional independence one, our test remains robust to the selection of dependence measures. We illustrate this by comparing the results of the proposed method using distance correlation (Székely et al., 2007) and improved projection correlation (Zhang and Zhu, 2024) in Appendix A.1.
We present an application of the proposed method to the problem of sufficient dimension reduction, specifically in testing whether an empirically estimated data representation satisfies the sufficiency requirement, as discussed in Section 4.
Our approach utilizes a permutation-based method for computing $p$-values. This method is supported by both simulation studies and real-data analysis, which collectively demonstrate its robustness, empirical validity, and practical utility. Theoretical justification for the permutation-based method is provided in Theorem 6, with numerical evaluations detailed in Section 4.
The remainder of the paper is organized as follows: We introduce our methodology in Section 2 and provide theoretical justifications in Section 3. Simulation studies and an application to dimension reduction are presented in Section 4. A brief conclusion is provided in Section 5. Reproducible code and implementation details are available at https://github.com/chenxuan-he/FlowCIT.
Notation. Throughout this paper, uppercase letters such as $X$, $Y$, and $Z$ denote random variables, with their realizations represented by the corresponding lowercase letters $x$, $y$, and $z$. Bold uppercase letters such as $\mathbf{X}$, $\mathbf{Y}$, and $\mathbf{Z}$ denote random vectors, while bold lowercase letters such as $\mathbf{x}$, $\mathbf{y}$, and $\mathbf{z}$ denote their realizations as real-valued vectors. The Euclidean norm of a vector $\mathbf{x}\in\mathbb{R}^{d}$ is denoted by $\|\mathbf{x}\|$. The Gamma function is denoted by $\Gamma(\cdot)$, the $d$-dimensional identity matrix by $\mathbf{I}_{d}$, and the indicator function by $\bm{1}(\cdot)$. We use =$\stackrel{{\scriptstyle d}}{{=}}$ to denote equal in distribution and $\stackrel{{\scriptstyle p}}{{\to}}$ to denote convergence in probability. The support of a random vector is denoted by $\textup{supp}(\cdot)$. For two sequences $a_{n}$ and $b_{n}$, the notation $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$ indicates that there exists a constant $C_{0}$ such that $|a_{n}|\leq C_{0}|b_{n}|$ as $n$ goes to infinity.
2 Methodology
In this section, we introduce our proposed method. The central idea is to utilize a transport map to convert a conditionally independent test problem into an unconditional one
2.1 From Conditional Independence to Independence
To begin with, we establish an equivalence between conditional independence and unconditional independence by invertible transport maps. Recall the formulation for conditional independence testing given in (1.1). Denote the original jointly distribution triplet as $(\mathbf{X},\mathbf{Y},\mathbf{Z})\in\mathbb{R}^{d_{\mathbf{x}}+d_{\mathbf{y}}+d_{\mathbf{z}}}$. Suppose there are two invertible transport maps that push the conditional distributions $\mathbf{X}\mid\mathbf{Z}$ and $\mathbf{Y}\mid\mathbf{Z}$ to two simple reference distributions (e.g., standard normal distributions). We say a function $f(\mathbf{x},\mathbf{z}):\mathbb{R}^{d_{\mathbf{x}}}\times\mathbb{R}^{d_{\mathbf{z}}}\to\mathbb{R}^{d_{\mathbf{x}}}$ is invertible with respect to the first argument if, for every fixed $\mathbf{z}\in\mathbb{R}^{d_{\mathbf{z}}}$, there exists an inverse function $f^{-1}:\mathbb{R}^{d_{\mathbf{x}}}\times\mathbb{R}^{d_{\mathbf{z}}}\to\mathbb{R}^{d_{\mathbf{x}}}$ such that $f^{-1}(f(\mathbf{x},\mathbf{z}),\mathbf{z})=f(f^{-1}(\mathbf{x},\mathbf{z}),\mathbf{z})=\mathbf{x}$ for all $\mathbf{x}\in\mathbb{R}^{d_{\mathbf{x}}}$ and $\mathbf{z}\in\mathbb{R}^{d_{\mathbf{z}}}$. We have the following Lemma 1.
Lemma 1 (Conditional independence to unconditional independence).
For the jointly distributed triplet $(\mathbf{X},\mathbf{Y},\mathbf{Z})\in\mathbb{R}^{d_{\mathbf{x}}+d_{\mathbf{y}}+d_{\mathbf{z}}}$, suppose there exist two invertible transport maps (with respect to their first arguments) $\bm{\xi}=f(\mathbf{X},\mathbf{Z})$ and $\bm{\eta}=g(\mathbf{Y},\mathbf{Z})$ such that:
$\bm{\xi}=$ $f(\mathbf{X},\mathbf{Z})\stackrel{{\scriptstyle d}}{{=}}\big{[}f(\mathbf{X},\mathbf{Z})\mid\mathbf{Z}=\mathbf{z}\big{]},\quad\forall\mathbf{z}\in\textup{supp}(\mathbf{Z}),$
$\bm{\eta}=$ $g(\mathbf{Y},\mathbf{Z})\stackrel{{\scriptstyle d}}{{=}}\big{[}g(\mathbf{Y},\mathbf{Z})\mid\mathbf{Z}=\mathbf{z}\big{]},\quad\forall\mathbf{z}\in\textup{supp}(\mathbf{Z}).$
Then, the following holds:
$\bm{\xi}\perp\!\!\!\perp\mathbf{Z},\quad\bm{\eta}\perp\!\!\!\perp\mathbf{Z}.$
Moreover, conditional independence in the original space is equivalent to independence in the transformed space:
$\mathbf{X}\perp\!\!\!\perp\mathbf{Y}\mid\mathbf{Z}\iff\bm{\xi}\perp\!\!\!\perp\bm{\eta}.$
By applying Lemma 1, we can convert the conditional independence test into an unconditional independence test. This transformation simplifies the analysis and allows us to leverage classical statistical measures of independence.
Moreover, if the transformed variables $\bm{\xi}$ and $\bm{\eta}$ follow normal distributions, classical measures of independence become well-defined and applicable. For example, Definition 2 (Székely et al., 2007) of distance correlation requires the existence of finite first-moments, a condition that is naturally satisfied by normal distributions. Detailed proofs of Lemma 1 can be found in Appendix B.
Lemma 1 establishes a practical framework for conducting conditional independence testing. The conditional Continuous Normalizing Flow model (CNF, Liu et al. 2023, Albergo and Vanden-Eijnden 2022, Lipman et al. 2022), a state-of-the-art generative model, is able to formulate the transport map and is proven to be well-posed under certain regularity conditions (Gao et al., 2024a). Building on this, the next section presents an implementation of Lemma 1 using conditional CNFs.
2.2 Conditional Continuous Normalizing Flows
In this section, we introduce the conditional CNFs with flow matching (Liu et al., 2023; Albergo and Vanden-Eijnden, 2022; Lipman et al., 2022). Let $\mathbf{X}_{0}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{x}}})$, $\mathbf{Y}_{0}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{y}}})$ be two random vectors independent of $(\mathbf{X},\mathbf{Y},\mathbf{Z})$. We illustrate the conditional CNF for the pair of distributions, the distribution of $\mathbf{X}_{0}$ and the conditional distribution of $(\mathbf{X}\mid\mathbf{Z})$; and similarly for $\mathbf{Y}_{0}$ and $(\mathbf{Y}\mid\mathbf{Z})$. A conditional rectified flow is constructed based on a linear stochastic interpolation between $\mathbf{X}_{0}$ and $\mathbf{X}$ over time $t\in[0,1]$:
$\mathbf{X}_{t}=(1-t)\mathbf{X}_{0}+t\mathbf{X},\quad\forall t\in[0,1].$ (2.1)
When $t=0$, since $\mathbf{X}_{0}\perp\!\!\!\perp(\mathbf{X},\mathbf{Z})$, it follows that $(\mathbf{X}_{0}\mid\mathbf{Z})=\mathbf{X}_{0}$ in distribution. When $t=1$, we have $(\mathbf{X}\mid\mathbf{Z})=(\mathbf{X}_{1}\mid\mathbf{Z})$. The process defined in (2.1) indicates that as $t$ goes from 0 to 1, the conditional distribution of $(\mathbf{X}_{t}\mid\mathbf{Z})$ progressively evolves from the standard normal distribution to the conditional distribution of $(\mathbf{X}_{1}\mid\mathbf{Z}),t\in[0,1].$
We now present the definition of the velocity field related to the conditional distribution of $(\mathbf{X}_{t}\mid\mathbf{Z})$.
Definition 1 (Velocity field).
For the time-differentiable random process $\{(\mathbf{X}_{t}\mid\mathbf{Z}),t\in[0,1]\}$ defined in (2.1), its velocity field $\mathbf{v}_{\mathbf{x}}(t,\mathbf{x},\mathbf{z})$ is defined as
$\mathbf{v}_{\mathbf{x}}(t,\mathbf{x},\mathbf{z})=$ $\mathbb{E}\left\{\frac{\textup{d}\mathbf{X}_{t}}{\textup{d}t}\ \bigg{|}\ \mathbf{X}_{t}=\mathbf{x},\mathbf{Z}=\mathbf{z}\right\}$
$=$ $\mathbb{E}\left(\mathbf{X}-\mathbf{X}_{0}\mid\mathbf{X}_{t}=\mathbf{x},\mathbf{Z}=\mathbf{z}\right),\quad\forall\mathbf{x}\in\textup{supp}(\mathbf{X}_{t}).$
By integrating the velocity field starting from $\mathbf{X}_{0}$, we derive the forward process:
$(\mathbf{X}_{t}^{\sharp}\mid\mathbf{Z})=\mathbf{X}_{0}+\int_{0}^{t}\mathbf{v}_{\mathbf{x}}(t,\mathbf{X}_{t}^{\sharp},\mathbf{Z})\textup{d}t,\quad\forall t\in[0,1],\quad\mathbf{X}^{\sharp}_{0}=\mathbf{X}_{0},$ (2.2)
where we use the notation $(\mathbf{X}_{t}^{\sharp}\mid\mathbf{Z})$ to represent the state of the process at time $t$ conditional on $\mathbf{Z}$. We will use this notation repeatedly below. Conversely, by integrating backward from $\mathbf{X}_{1}=\mathbf{X},$ we obtain the reverse process:
$(\mathbf{X}_{t}^{\dagger}\mid\mathbf{Z})=\mathbf{X}_{1}+\int_{1}^{t}\mathbf{v}_{\mathbf{x}}(t,\mathbf{X}_{t}^{\dagger},\mathbf{Z})\textup{d}t,\quad\forall t\in[0,1],\quad(\mathbf{X}^{\dagger}_{1}\mid\mathbf{Z})=(\mathbf{X}\mid\mathbf{Z}).$ (2.3)
The forward process $(\mathbf{X}_{t}^{\sharp}\mid\mathbf{Z})$ and the reverse process $(\mathbf{X}_{t}^{\dagger}\mid\mathbf{Z})$ have been shown to possess the marginal preserving property, as detailed in Lemma 2 below.
Lemma 2 (Marginal preserving property).
Under several regularity conditions, these two processes $(\mathbf{X}_{t}^{\sharp}\mid\mathbf{Z})$ and $(\mathbf{X}_{t}^{\dagger}\mid\mathbf{Z})$ defined in (2.2) and (2.3) are well-posed and unique. In addition, the following property holds:
$(\mathbf{X}_{t}\mid\mathbf{Z}=\mathbf{z})\stackrel{{\scriptstyle d}}{{=}}(\mathbf{X}_{t}^{\sharp}\mid\mathbf{Z}=\mathbf{z})\stackrel{{\scriptstyle d}}{{=}}(\mathbf{X}_{t}^{\dagger}\mid\mathbf{Z}=\mathbf{z}),\quad\text{for any }(\mathbf{z},t)\in\textup{supp}(\mathbf{Z})\times[0,1],$
where $(\mathbf{X}_{t}\mid\mathbf{Z})$ is defined in (2.1).
The marginal preserving property is well-established in the literature. For instance, Theorem 5.1 and Corollary 5.2 in Gao et al. (2024a) establish this property under suitable assumptions. Similar results can also be found in Liu (2022), Albergo et al. (2023), and Huang et al. (2024a).
According to Lemma 2, the velocity field $\mathbf{v}_{\mathbf{x}}(t,\mathbf{x},\mathbf{z})$ enables the transformation of the distribution $\mathbf{X}_{0}$ into $(\mathbf{X}\mid\mathbf{Z})$, and vice versa. Furthermore, since the ordinary differential equation (ODE) in (2.1) and the integration process defined in (2.3) are deterministic, the transformation is one-to-one and invertible. This property is not shared by the stochastic differential equation-based methods, such as diffusion models (Ho et al., 2020; Song et al., 2020). Consequently, the transport map induced by the rectified flow is invertible and meets the conditions outlined in Lemma 1. Thus, we introduce two rectified flows:
$\mathbf{X}_{t}=$ $(1-t)\mathbf{X}_{0}+t\mathbf{X},\quad\forall t\in[0,1],$
$\mathbf{Y}_{t}=$ $(1-t)\mathbf{Y}_{0}+t\mathbf{Y},\quad\forall t\in[0,1].$
These two processes induce two velocity fields $\mathbf{v}_{\mathbf{x}}(t,\mathbf{x},\mathbf{z})$ and $\mathbf{v}_{\mathbf{y}}(t,\mathbf{y},\mathbf{z})$ as given in Definition 1. At the population level, by integrating the velocity fields starting from $\mathbf{X}$ and $\mathbf{Y},$ we obtain:
$(\mathbf{X}_{0}^{\dagger}\mid\mathbf{Z})$ $=\mathbf{X}+\int_{1}^{0}\mathbf{v}_{\mathbf{x}}(t,\mathbf{X}_{t}^{\dagger},\mathbf{Z})\textup{d}t,$
$(\mathbf{Y}_{0}^{\dagger}\mid\mathbf{Z})$ $=\mathbf{Y}+\int_{1}^{0}\mathbf{v}_{\mathbf{y}}(t,\mathbf{Y}_{t}^{\dagger},\mathbf{Z})\textup{d}t.$
By Lemma 2, we have $(\mathbf{X}_{0}^{\dagger}\mid\mathbf{Z})\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}_{0}$ and $(\mathbf{Y}_{0}^{\dagger}\mid\mathbf{Z})\stackrel{{\scriptstyle d}}{{=}}\mathbf{Y}_{0}$. We define:
$\bm{\xi}:=$ $(\mathbf{X}_{0}^{\dagger}\mid\mathbf{Z})\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}_{0},$ (2.4)
$\bm{\eta}:=$ $(\mathbf{Y}_{0}^{\dagger}\mid\mathbf{Z})\stackrel{{\scriptstyle d}}{{=}}\mathbf{Y}_{0},$ (2.5)
where $\bm{\xi}$ is a function of $\mathbf{X}$ and $\mathbf{Z}$, and $\bm{\eta}$ is a function of $\mathbf{Y}$ and $\mathbf{Z}$. The pair $(\bm{\xi},\bm{\eta})$ is exactly what we expect in Lemma 1. We illustrate this idea in Example 1 below.
Example 1.
We illustrate Lemma 1 and Lemma 2 using a Gaussian example with a univariate triplet $(X,Y,Z)$. Suppose $Z\sim\text{Normal}(0,1)$, $X\mid Z\sim\text{Normal}(Z,1)$, and $Y\mid Z\sim\text{Normal}(Z,1)$. We know that $X\perp\!\!\!\perp Y\mid Z$. The linear interpolation between $X_{0}$ and $X$ is
$X_{t}=(1-t)X_{0}+tX,\quad\forall t\in[0,1].$
The joint distribution of $\left((X,X_{0},X_{t})^{\top}\mid Z\right)$ is $\text{Normal}\left(\bm{\mu},\Sigma\right)$, where $\bm{\mu}=(Z,0,Z-tZ)^{\top}$ and
$\Sigma=\left(\begin{array}[]{ccc}1&0&t\\ 0&1&1-t\\ t&1-t&1-2t-t^{2}\end{array}\right).$
By the property of joint normal distribution, the velocity field is given by
$v_{x}(t,x,z)=\mathbb{E}(X-X_{0}\mid X_{t}=x,Z=z)=\frac{2t-1}{1-2t-2t^{2}}\{x-(1-t)z\}=\frac{\textup{d}x}{\textup{d}t}.$
Solving the ODE above, with initial value $x$ at $t=1$, the value at $t=0$ is $z-x$. Thus, we have $\xi=Z-X\sim\text{Normal}(0,1)$. Similarly, we have $\eta=Z-Y\sim\text{Normal}(0,1)$. Thus, $\xi\perp\!\!\!\perp\eta$ is satisfied.
2.3 Independence Measures
By transforming the conditional variables $(\mathbf{X}\mid\mathbf{Z})$ and $(\mathbf{Y}\mid\mathbf{Z})$ into $\bm{\xi}$ and $\bm{\eta}$, respectively, the conditional independence testing problem $\mathbf{X}\perp\!\!\!\perp\mathbf{Y}\mid\mathbf{Z}$ is simplified to $\bm{\xi}\perp\!\!\!\perp\bm{\eta}$, as stated in Lemma 1. Subsequently, we use independence measures to evaluate the dependence between $\bm{\xi}$ and $\bm{\eta}$.
Since $\bm{\xi}$ and $\bm{\eta}$ are Gaussian-distributed—a well-behaved distribution with finite moments—several popular correlation measures are suitable for use. These include Distance Correlation (DC, Székely et al. 2007), Projection Correlation (PC, Zhu et al. 2017), and Improved Projection Correlation (IPC, Zhang and Zhu 2024). They are all equal to zero if and only if $\bm{\xi}$ and $\bm{\eta}$ are independent.
Given the computational efficiency of DC and IPC, we recommend using these two measures to assess the dependence of $\bm{\xi}$ and $\bm{\eta}$. Our numerical studies suggest that the test results remain robust regardless of the chosen dependence measure. For further details, please refer to Appendix A.1. In the following, we illustrate the idea of independence measures by distance covariance and distance correlation. An introduction for IPC is given in Appendix A.1.1.
The distance covariance between two random vectors $\bm{\xi}\in\mathbb{R}^{d_{\mathbf{x}}}$ and $\bm{\eta}\in\mathbb{R}^{d_{\mathbf{y}}}$ is defined by
$\textup{dcov}^{2}(\bm{\xi},\bm{\eta})=$ $\int_{\mathbb{R}^{d_{\mathbf{x}}+d_{\mathbf{y}}}}\left\|f_{\bm{\xi},\bm{\eta}}(\mathbf{u},\mathbf{v})-f_{\bm{\xi}}(\mathbf{u})f_{\bm{\eta}}(\mathbf{v})\right\|^{2}w(\mathbf{u},\mathbf{v})\textup{d}\mathbf{u}\textup{d}\mathbf{v},$ (2.6)
where $f_{\bm{\xi},\bm{\eta}}(\cdot,\cdot)$ is the joint characteristic function of $(\bm{\xi},\bm{\eta})$, $f_{\bm{\xi}}(\cdot)$ and $f_{\bm{\eta}}(\cdot)$ are the marginal characteristic functions of $\bm{\xi}$ and $\bm{\eta}$, and $w(\cdot,\cdot)$ is the weight function defined by $w(\mathbf{u},\mathbf{v})=\left\{c_{d_{\mathbf{x}}}c_{d_{\mathbf{y}}}\|\mathbf{u}\|^{1+d_{\mathbf{x}}}\|\mathbf{v}\|^{1+d_{\mathbf{y}}}\right\}^{-1},$ with $c_{d}=\pi^{(1+d)/2}/\Gamma((1+d)/2)$.
Since $\bm{\xi}\perp\!\!\!\perp\bm{\eta}\iff f_{\bm{\xi},\bm{\eta}}(\mathbf{u},\mathbf{v})=f_{\bm{\xi}}(\mathbf{u})f_{\bm{\eta}}(\mathbf{v})$, the distance covariance is nonnegative and equals to zero if and only if $\bm{\xi}$ and $\bm{\eta}$ are independent. According to Remark 3 in Székely et al. (2007), let $(\bm{\xi}_{1},\bm{\eta}_{1})$, $(\bm{\xi}_{2},\bm{\eta}_{2})$, and $(\bm{\xi}_{3},\bm{\eta}_{3})$ be independent copies of $(\bm{\xi},\bm{\eta})$. If $\mathbb{E}\|\bm{\xi}\|^{2}<\infty$ and $\mathbb{E}\|\bm{\eta}\|^{2}<\infty$, the distance covariance defined in (2.6) can be formulated as
$\textup{dcov}^{2}(\bm{\xi},\bm{\eta})=$ $\mathbb{E}\left(\|\bm{\xi}_{1}-\bm{\xi}_{2}\|\|\bm{\eta}_{1}-\bm{\eta}_{2}\|\right)+\mathbb{E}\|\bm{\xi}_{1}-\bm{\xi}_{2}\|\mathbb{E}\|\bm{\eta}_{1}-\bm{\eta}_{2}\|$
$-2\mathbb{E}\left(\|\bm{\xi}_{1}-\bm{\xi}_{2}\|\|\bm{\eta}_{1}-\bm{\eta}_{3}\|\right).$ (2.7)
This formulation allows us to more effectively provide the empirical version of distance covariance. The distance correlation is further defined by:
$\textup{dcorr}^{2}(\bm{\xi},\bm{\eta})=\frac{\textup{dcov}^{2}(\bm{\xi},\bm{\eta})}{\textup{dcov}(\bm{\xi},\bm{\xi})\textup{dcov}(\bm{\eta},\bm{\eta})}.$
2.4 Test Statistic
Based on the preceding discussions, for the triplet $(\mathbf{X},\mathbf{Y},\mathbf{Z})$, we propose transforming the conditional distributions $(\mathbf{X}\mid\mathbf{Z})$ and $(\mathbf{Y}\mid\mathbf{Z})$ into $\bm{\xi}$ and $\bm{\eta}$ by conditional CNFs. The dependence between the transformed variables $\bm{\xi}$ and $\bm{\eta}$ is then assessed via independence measures. In this section, we develop the empirical counterpart of this procedure to perform the conditional independence test.
Consider the observed dataset $\mathcal{D}_{n}=\{(\mathbf{X}_{i},\mathbf{Y}_{i},\mathbf{Z}_{i})\}_{i=1}^{n}$. The proposed test statistic is constructed through four key steps. First, we estimate velocity fields via deep neural networks. Next, we compute the transformed variables $\bm{\xi}$ and $\bm{\eta}$ using the maps defined in (2.4) and (2.5). Third, we construct the test statistic $T_{n}$ as the sample version of the chosen independence measure. Finally, we compute the $p$-value using permutation-based methods. We now give detailed explanations of the four steps.
Step 1. Estimation of the velocity fields $\mathbf{v}_{\mathbf{x}}(t,\mathbf{x},\mathbf{z})$ and $\mathbf{v}_{\mathbf{y}}(t,\mathbf{y},\mathbf{z})$.
In the first step, we estimate the velocity fields using deep neural networks. As established in Definition 1, these velocity fields correspond to conditional expectations, making the least squares estimator a natural choice for their estimation. Operationally, we augment the observed dataset $\mathcal{D}_{n}=\{(\mathbf{X}_{i},\mathbf{Y}_{i},\mathbf{Z}_{i})\}_{i=1}^{n}$ with three auxiliary random samples: $\{\mathbf{X}_{0,i}\}_{i=1}^{n}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{x}}})$, $\{\mathbf{Y}_{0,i}\}_{i=1}^{n}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{y}}})$, and $\{t_{i}\}_{i=1}^{n}\sim\text{Uniform}([0,1])$. The velocity field estimators are obtained by minimizing the empirical squared losses:
$\hat{\mathbf{v}}_{\mathbf{x}}(t,\mathbf{x},\mathbf{z})$ $=\operatorname*{arg\,min}_{\mathbf{v}_{\mathbf{x}}\in\mathcal{F}_{n,\mathbf{x}}}\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{v}_{\mathbf{x}}(t_{i},\mathbf{X}_{i},\mathbf{Z}_{i})-(\mathbf{X}_{i}-\mathbf{X}_{0,i})\|_{2}^{2},$ (2.8)
$\hat{\mathbf{v}}_{\mathbf{y}}(t,\mathbf{y},\mathbf{z})$ $=\operatorname*{arg\,min}_{\mathbf{v}_{\mathbf{y}}\in\mathcal{F}_{n,\mathbf{y}}}\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{v}_{\mathbf{y}}(t_{i},\mathbf{Y}_{i},\mathbf{Z}_{i})-(\mathbf{Y}_{i}-\mathbf{Y}_{0,i})\|_{2}^{2},$ (2.9)
where $\mathcal{F}_{n,\mathbf{x}}$ and $\mathcal{F}_{n,\mathbf{y}}$ denote neural network function classes (see Section 3.1 for details). These estimated velocity fields then serve as the foundation for computing the transformed variables $\bm{\xi}$ and $\bm{\eta}$ in the subsequent step.
Step 2. Estimation of $\bm{\xi}$ and $\bm{\eta}$.
At the population level, by Lemmas 12, the dataset $\mathcal{D}_{n}=\{(\mathbf{X}_{i},\mathbf{Y}_{i},\mathbf{Z}_{i})\}_{i=1}^{n}$ corresponds to a transformed dataset $\mathcal{D}_{n}^{\prime}=\{(\bm{\xi}_{i},\bm{\eta}_{i})\}_{i=1}^{n}$. In this step, we obtain an estimate of $\mathcal{D}_{n}^{\prime}$, denoted by $\mathcal{D}_{n}^{\prime\prime}=\{(\hat{\bm{\xi}}_{i},\hat{\bm{\eta}}_{i})\}_{i=1}^{n}$:
$\hat{\bm{\xi}}_{i}$ $=\mathbf{X}_{i}+\int_{1}^{0}\hat{\mathbf{v}}_{\mathbf{x}}(t,\mathbf{X}_{t,i}^{\dagger},\mathbf{Z}_{i})\textup{d}t,\quad i=1,\ldots,n,$ (2.10)
$\hat{\bm{\eta}}_{i}$ $=\mathbf{Y}_{i}+\int_{1}^{0}\hat{\mathbf{v}}_{\mathbf{y}}(t,\mathbf{X}_{t,i}^{\dagger},\mathbf{Z}_{i})\textup{d}t,\quad i=1,\ldots,n.$ (2.11)
After obtaining $\mathcal{D}_{n}^{\prime\prime}=\{(\hat{\bm{\xi}}_{i},\hat{\bm{\eta}}_{i})\}_{i=1}^{n}$, we proceed to the next step to construct the test statistic.
Step 3. Construction of the test statistic by $\hat{\bm{\xi}}$ and $\hat{\bm{\eta}}$.
We compute the sample version of the chosen independence measure based on $\mathcal{D}_{n}^{\prime\prime}=\{(\hat{\bm{\xi}}_{i},\hat{\bm{\eta}}_{i})\}_{i=1}^{n}$ and take it as the test statistic. For instance, if we choose distance correlation, we first compute the sample version of distance covariance:
$\textup{dcov}_{n}^{2}(\hat{\bm{\xi}},\hat{\bm{\eta}})=S_{1}+S_{2}-2S_{3},$
where
$S_{1}=$ $\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{l}\|,$
$S_{2}=$ $\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{l}\|,$
$S_{3}=$ $\frac{1}{n^{3}}\sum_{k=1}^{n}\sum_{l,m=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{m}\|.$
The estimated distance correlation is used as the test statistic and given by:
$T_{n}=\textup{dcorr}_{n}^{2}(\hat{\bm{\xi}},\hat{\bm{\eta}})=\frac{\textup{dcov}_{n}^{2}(\hat{\bm{\xi}},\hat{\bm{\eta}})}{\textup{dcov}_{n}(\hat{\bm{\xi}},\hat{\bm{\xi}})\textup{dcov}_{n}(\hat{\bm{\eta}},\hat{\bm{\eta}})},$ (2.12)
where $\textup{dcov}_{n}(\hat{\bm{\xi}},\hat{\bm{\xi}})$ and $\textup{dcov}_{n}(\hat{\bm{\eta}},\hat{\bm{\eta}})$ are computed in a manner similar to $\textup{dcov}_{n}(\hat{\bm{\xi}},\hat{\bm{\eta}})$. After obtaining the test statistic, we compute the $p$-value by the permutation-based method.
Step 4. Obtaining the $p$-value by permutation.
To get a $p$-value for the test statistic $T_{n}$, we first randomly permute $\mathcal{D}_{n}^{\prime\prime}$ as $\mathcal{D}_{n,b}^{\prime\prime}=\{(\hat{\bm{\xi}}_{i},\hat{\bm{\eta}}_{\pi_{b}(i)})\}_{i=1}^{n}$, where $\pi_{b}$ is a uniformly random permutation of $\{1,2,\cdots,n\}$. A permuted statistic $T_{n,b}$ is then computed similar to (2.12) based on $\mathcal{D}_{n,b}^{\prime\prime}$. We replicate the permutation $B$ times and compute the $p$-value as following:
$p=\frac{1}{B}\sum_{b=1}^{B}\bm{1}(T_{n,b}\geq T_{n}).$ (2.13)
We summarize the full procedure in Algorithm 1 below.
Algorithm 1 FlowCIT: Conditional independence test based on CNFs.
1:Dataset $\mathcal{D}_{n}=\{(\mathbf{X}_{i},\mathbf{Y}_{i},\mathbf{Z}_{i})\}_{i=1}^{n}$; Number of permutations $B$.
2:Test statistic $T_{n}$; $p$-value.
3:Generate $\{\mathbf{X}_{0,i}\}_{i=1}^{n}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{x}}})$, $\{\mathbf{Y}_{0,i}\}_{i=1}^{n}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{y}}})$, and $\{t_{i}\}_{i=1}^{n}\sim\text{Uniform}([0,1])$.
4:(Step 1) Estimate the velocity fields $\mathbf{v}_{\mathbf{x}}$ and $\mathbf{v}_{\mathbf{y}}$ by $\{(\mathbf{X}_{i},\mathbf{Y}_{i},\mathbf{Z}_{i})\}_{i=1}^{n}$, $\{\mathbf{X}_{0,i}\}_{i=1}^{n}$, $\{\mathbf{Y}_{0,i}\}_{i=1}^{n}$, $\{t_{i}\}_{i=1}^{n}$ as given in (2.8) and (2.9).
5:(Step 2) Compute $\hat{\bm{\xi}}_{i}$ and $\hat{\bm{\eta}}_{i}$ based on $\hat{\mathbf{v}}_{\mathbf{x}}$ and $\hat{\mathbf{v}}_{\mathbf{y}}$ as given in (2.10) and (2.11). Denote the estimated version by $\mathcal{D}_{n}^{\prime\prime}=\{(\hat{\bm{\xi}}_{i},\hat{\bm{\eta}}_{i})\}_{i=1}^{n}$.
6:(Step 3) Compute the test statistic $T_{n}$ by an independence measure based on $\mathcal{D}_{n}^{\prime\prime}$ (e.g., distance correlation by (2.12)).
7:(Step 4) For $b=1,\dots,B$ do
  • Generate a uniformly random permutation $\pi_{b}$ of $\{1,2,\dots,n\}$.
  • Denote the permuted dataset as $\mathcal{D}_{n,b}^{\prime\prime}=\{(\hat{\bm{\xi}}_{i},\hat{\bm{\eta}}_{\pi_{b}(i)})\}_{i=1}^{n}$.
  • Compute the permuted test statistic $T_{n,b}$ based on $\mathcal{D}_{n,b}^{\prime\prime}$.
  • end for
8:Calculate the $p$-value as given in (2.13).
9:return $T_{n}$; $p$-value.
3 Theoretical Analysis
In this section, we provide a theoretical analysis of the proposed test. As outlined in Algorithm 1, our procedure involves a four-step estimation process: estimating the velocity field, estimating the transformed pairs, calculating the dependence measure of the transformed pairs, and estimating the $p$-value.
To facilitate the analysis, we begin by introducing the deep neural networks in Section 3.1. We then give the consistency of the estimated velocity field by deep neural networks and the consistency of the estimated transport maps in Lemma 4. Furthermore, we prove the consistency of the test statistic $T_{n}$ and the $p$-value obtained via permutation in Theorems 56.
3.1 Deep Neural Networks
We provide a brief review of regression function estimators based on feedforward neural networks with the rectified linear unit (ReLU) activation function, defined as $\sigma(x):=\max(0,x)$. The function class $\mathcal{F}_{n}$ is specified as $\mathcal{F}_{\mathcal{T},\mathcal{W},\mathcal{U},\mathcal{S},\mathcal{B}}$, representing a class of feedforward neural networks $f_{\phi}:\mathbb{R}^{d}\to\mathbb{R}$ with parameters $\phi$, depth $\mathcal{T}$, width $\mathcal{W}$, size $\mathcal{S}$, and number of neurons $\mathcal{U}$. Moreover, each network satisfies $\|f_{\phi}\|_{\infty}\leq\mathcal{B}$ for some $0<\mathcal{B}<\infty$. Note that the parameters may be indexed by $n$, and we omit $n$ for simplicity.
To be specific, let $\mathcal{T}$ denote the depth of the hidden layers, and the total number of layers is $\mathcal{T}+2$. The width of each layer is represented by the $(\mathcal{T}+2)$-vector $(p_{0},\ldots,p_{\mathcal{T}+1})^{\top}$, where $\mathcal{W}$ is defined as $\mathcal{W}:=\max(p_{1},\ldots,p_{\mathcal{T}})$. The total number of neurons $\mathcal{U}$ is defined as $\mathcal{U}=\sum_{i=1}^{\mathcal{T}}p_{i}$. Additionally, $\mathcal{S}$ represents the total number of computational units or parameters, satisfying $\mathcal{S}=\sum_{i=0}^{\mathcal{T}}\{p_{i+1}(p_{i}+1)\}=O(\mathcal{W}^{2}\mathcal{T})$.
3.2 Consistency of the Test Statistic
In this section, we provide the theoretical properties of our test. First, we give the convergence of the estimated velocity fields $\hat{\mathbf{v}}_{\mathbf{x}}$ and $\hat{\mathbf{v}}_{\mathbf{y}}$ by Lemma 3, a result that has already been established in the literature (Gao et al., 2024b).
Lemma 3 (Convergence of flow matching).
Suppose that Lemma 2 holds for $\mathbf{X}_{t}^{\dagger}\mid\mathbf{Z}$ and $\mathbf{Y}_{t}^{\dagger}\mid\mathbf{Z}$, and the velocity fields $\mathbf{v}_{\mathbf{x}}(t,\mathbf{x},\mathbf{z})$ and $\mathbf{v}_{\mathbf{y}}(t,\mathbf{y},\mathbf{z})$ have the following regularity properties:
  1. 1.
    For any $s,t\in[0,1-\underline{t}]$, $\mathbf{x}\in\mathbb{R}^{d_{\mathbf{x}}}$, $\mathbf{y}\in\mathbb{R}^{d_{\mathbf{y}}}$, and $\mathbf{z}\in\mathbb{R}^{d_{\mathbf{z}}}$, $\|\mathbf{v}_{\mathbf{x}}(t,\mathbf{x},\mathbf{z})-\mathbf{v}_{\mathbf{x}}(s,\mathbf{x},\mathbf{z})\|\leq L_{t}|t-s|$ and $\|\mathbf{v}_{\mathbf{y}}(t,\mathbf{y},\mathbf{z})-\mathbf{v}_{\mathbf{y}}(s,\mathbf{y},\mathbf{z})\|\leq L_{t}|t-s|$ with $L_{t}\lesssim\underline{t}^{-2}$;
  2. 2.
    For any $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathbb{R}^{d_{\mathbf{x}}}$, $\mathbf{y}_{1},\mathbf{y}_{2}\in\mathbb{R}^{d_{\mathbf{y}}}$, $\mathbf{z}_{1},\mathbf{z}_{2}\in\mathbb{R}^{d_{\mathbf{z}}}$, and $t\in[0,1]$, $\|\mathbf{v}_{\mathbf{x}}(t,\mathbf{x}_{1},\mathbf{z}_{1})-\mathbf{v}_{\mathbf{x}}(t,\mathbf{x}_{2},\mathbf{z}_{2})\|\leq L_{\mathbf{z}}\|(\mathbf{x}_{1},\mathbf{z}_{1})-(\mathbf{x}_{2},\mathbf{z}_{2})\|$ and $\|\mathbf{v}_{\mathbf{y}}(t,\mathbf{y}_{1},\mathbf{z}_{1})-\mathbf{v}_{\mathbf{y}}(t,\mathbf{y}_{2},\mathbf{z}_{2})\|\leq L_{\mathbf{z}}\|(\mathbf{y}_{1},\mathbf{z}_{1})-(\mathbf{y}_{2},\mathbf{z}_{2})\|$ with $L_{\mathbf{z}}\lesssim 1$.
Then, with the width $\mathcal{W}$, size $\mathcal{S}$, and depth $\mathcal{T}$ of the neural network function classes $\mathcal{F}_{n,\mathbf{x}}$ and $\mathcal{F}_{n,\mathbf{y}}$ properly specified, the flow matching errors converge to zero in probability as $n\to\infty$:
$\mathbb{E}_{(t,\mathbf{X},\mathbf{Z})}\left\|\hat{\mathbf{v}}_{\mathbf{x}}(t,\mathbf{X},\mathbf{Z})-\mathbf{v}_{\mathbf{x}}(t,\mathbf{X},\mathbf{Z})\right\|^{2}$ $\stackrel{{\scriptstyle p}}{{\to}}0,$
$\mathbb{E}_{(t,\mathbf{Y},\mathbf{Z})}\left\|\hat{\mathbf{v}}_{\mathbf{y}}(t,\mathbf{Y},\mathbf{Z})-\mathbf{v}_{\mathbf{y}}(t,\mathbf{Y},\mathbf{Z})\right\|^{2}$ $\stackrel{{\scriptstyle p}}{{\to}}0.$
Recent studies, including Gao et al. (2024b), Fukumizu et al. (2025), and Cai and Li (2025), have explored the error analysis of flow matching. We provide a detailed discussion of Lemma 3 in Appendix D.1. Building on Lemma 3, we introduce Lemma 4, which establishes the convergence of the estimated $\bm{\xi}$ and $\bm{\eta}$.
Lemma 4 (Convergence of $\bm{\xi}$ and $\bm{\eta}$).
Suppose that the conditions for Lemmas 23 are satisfied. With $n\to\infty$, it follows that:
$\mathbb{E}\big{\|}\hat{\bm{\xi}}-\bm{\xi}\big{\|}^{2}\stackrel{{\scriptstyle p}}{{\to}}0,\quad\mathbb{E}\big{\|}\hat{\bm{\eta}}-\bm{\eta}\big{\|}^{2}\stackrel{{\scriptstyle p}}{{\to}}0.$
Lemma 4 suggests that with $n\to\infty$, the difference between the estimated $\hat{\bm{\xi}}$ and the true $\bm{\xi}$ converges to zero in probability. The proof of Lemma 4 is given in Section D.2. We are now ready to derive the consistency of our test statistic $T_{n}$ defined in (2.12).
Under $H_{0}$, according to Lemma 1, we have $\textup{dcorr}(\bm{\xi},\bm{\eta})=0$. Under $H_{1}$, there exists a $\zeta>0$ such that $\textup{dcorr}(\bm{\xi},\bm{\eta})=\zeta$. Theorem 5 establishes the consistency of the test statistic $T_{n}$ defined in (2.12) under both $H_{0}$ and $H_{1}$.
Theorem 5.
Suppose that the conditions for Lemmas 24 are satisfied. Under $H_{0}$, the test statistic $T_{n}\stackrel{{\scriptstyle p}}{{\to}}0$. Under $H_{1}$, the test statistic $T_{n}\stackrel{{\scriptstyle p}}{{\to}}\zeta=\textup{dcorr}(\bm{\xi},\bm{\eta})>0$.
The proof of Theorem 5 is given in Section C.1. In addition, we give the following Theorem 6 to indicate the consistency of the permutation-based method given in Algorithm 1, with proof deferred to Section C.2.
Theorem 6.
Suppose that the conditions for Lemmas 24 are satisfied. Under $H_{0}$, $\mbox{Pr}(T_{n,b}\leq t)-\mbox{Pr}(T_{n}\leq t)\stackrel{{\scriptstyle p}}{{\to}}0$ for all $t\in\mathbb{R}^{+}$. Under $H_{1}$, $\mbox{Pr}(T_{n,b}\geq T_{n})\stackrel{{\scriptstyle p}}{{\to}}0$.
Theorems 56 provide the theoretical guarantees to perform the proposed test. In Section 4, we validate the robust performance of our proposed test through numerical studies.
4 Numerical Studies
In the numerical studies, we validate our proposed FlowCIT given in Algorithm 1 through simulations and real-data analysis. We compare our test with the following nonparametric conditional independence test methods: kernel-based conditional independence test (KCI, Zhang et al. 2011), conditional distance correlation test (CDC, Wang et al. 2015), classifier conditional independence test (CCIT, Sen et al. 2017), fast conditional independence test (FCIT, Chalupka et al. 2018), and a test proposed by Cai et al. (2022) (abbreviated as CLZ). The KCI test is implemented by the R package CondIndTests::KCI, the CDC and FCIT are implemented by the Python package hyppo (Panda et al., 2024), the CCIT is implemented by its Python package CCIT, and the CLZ test is implemented by its original R code. Please refer to Appendix A.3 for the details of implementations of these methods.
4.1 Simulations
We consider four different examples in Models 14, covering 16 settings. These models range from linear to nonlinear examples and include univariate, low-dimensional, and moderately high-dimensional scenarios.
As a prelude, we summarize the comparisons of these methods based on our simulation models in Table 1. Additionally, our method effectively controls the type-I error under $H_{0}$ and achieves the highest statistical power performance across all simulation models.
Table 1: Comparisons of nonparametric conditional independence test methods across our simulation settings. A checkmark ($\checkmark$) indicates that the method is workable under the given setting—effectively controlling the type-I error, demonstrating statistical power against alternatives, and being computationally feasible. A triangle ($\Delta$) indicates that the method is partially workable. Cells are left blank if a method fails to show detectable statistical power under our model settings or is computationally infeasible in practice.
Dimensions Examples FlowCIT KCI CDC FCIT CCIT CLZ
$\mathbf{Z}$ $\mathbf{X}$ and $\mathbf{Y}$
Low Univariate Model 1 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Low Model 2 $\checkmark$ $\checkmark$ $\checkmark$ $\Delta$
High Low Model 3 $\checkmark$ $\Delta$ $\Delta$
High Model 4 $\checkmark$ $\Delta$ $\Delta$
In these models, we use the parameter $\psi$ to control the deviation from $H_{0}$: if $\psi=0$, the model is under $H_{0}$; otherwise, a larger $|\psi|$ indicates a greater deviation. We use plots to visualize the empirical type-I errors and statistical powers across various settings. Since the data points are densely clustered under $H_{0}$, detailed empirical type-I errors are provided in Appendix A.2. In this section, our method is implemented using distance correlation, and a comparison between distance correlation and improved projection correlation is presented in Appendix A.1, which suggests that their performance is very similar. All results are based on 200 replications.
4.1.1 Low-dimensional $\mathbf{Z}$
We begin with two low-dimensional $\mathbf{Z}$ examples (Models 12). In Model 1, we consider a simple example where $X$ and $Y$ are univariate and $\mathbf{Z}$ is two-dimensional. Specifically, in Settings 1.31.4, we consider heavy-tailed distributions.
Model 1 (Univariate $X$ and $Y$, and low-dimensional $\mathbf{Z}$).
Let $(d_{\mathbf{x}},d_{\mathbf{y}},d_{\mathbf{z}})=(1,1,2)$ and the sample size $n=500$. Each element of $\mathbf{B}_{1}\in\mathbb{R}^{d_{\mathbf{z}}\times d_{\mathbf{x}}}$, $\mathbf{B}_{2}\in\mathbb{R}^{d_{\mathbf{z}}\times d_{\mathbf{y}}}$, $\mathbf{B}_{3}\in\mathbb{R}^{d_{\mathbf{x}}\times d_{\mathbf{y}}}$, and $\mathbf{Z}\in\mathbb{R}^{d_{\mathbf{z}}}$ are generated from $\text{Normal}(0,1)$. Variables $X$ and $Y$ are generated under Settings 1.11.4, and we vary $\psi$ over the set $\{0,0.1,0.2,0.3,0.4\}$. For the error terms, we generate $\epsilon_{\mathbf{y}}\sim\text{Normal}(0,1)$, and $\epsilon_{\mathbf{x}}$ is set differently. Let $t_{3}$ denote the $t$-distribution with 3 degrees of freedom.
Setting 1.1.
$X=\mathbf{Z}\mathbf{B}_{1}+\epsilon_{\mathbf{x}},\quad Y=\mathbf{Z}\mathbf{B}_{2}+\psi X\mathbf{B}_{3}+\epsilon_{\mathbf{y}},\quad\epsilon_{\mathbf{x}}\sim\text{Normal}(0,1)$.
Setting 1.2.
$X=\mathbf{Z}\mathbf{B}_{1}+\epsilon_{\mathbf{x}},\quad Y=\mathbf{Z}\mathbf{B}_{2}+\exp(\psi X\mathbf{B}_{3})+\epsilon_{\mathbf{y}},\quad\epsilon_{\mathbf{x}}\sim\text{Normal}(0,1)$.
Setting 1.3.
$X=\mathbf{Z}\mathbf{B}_{1}+\epsilon_{\mathbf{x}},\quad Y=\mathbf{Z}\mathbf{B}_{2}+\psi X\mathbf{B}_{3}+\epsilon_{\mathbf{y}},\quad\epsilon_{\mathbf{x}}\sim t_{3}$.
Setting 1.4.
$X=\cos(\mathbf{Z}\mathbf{B}_{1})+\epsilon_{\mathbf{x}},\quad Y=\mathbf{Z}\mathbf{B}_{2}+\psi X\mathbf{B}_{3}+\epsilon_{\mathbf{y}},\quad\epsilon_{\mathbf{x}}\sim t_{3}$.
The results of Model 1 are shown in Figure 1, where the $x$-axis represents $\psi$. When $\psi=0$, all six methods effectively control the type-I error under $H_{0}$. As $\psi$ increases, our proposed FlowCIT consistently achieves the highest statistical power under $H_{1}$. The second and the third most powerful methods are KCI and CDC tests, respectively. In contrast, CCIT is the most conservative method. In the next Model 2, we analyze multivariate pairs $\mathbf{X}$ and $\mathbf{Y}$, with all three vectors specified as three-dimensional.
Figure 1: Empirical type-I errors and statistical powers of Model 1 with sample size $n=500$. The dashed gray line below refers to significance level of $0.05$. The $x$-axis refers to the value of $\psi$.
Model 2 (Multivariate $\mathbf{X}$, $\mathbf{Y}$, and $\mathbf{Z}$).
Let $(d_{\mathbf{x}},d_{\mathbf{y}},d_{\mathbf{x}})=(3,3,3)$ and the sample size $n=500$. Each element of $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, $\mathbf{B}_{3}$, and $\mathbf{Z}$ are generated from $\text{Normal}(0,1)$. Vectors $\mathbf{X}$ and $\mathbf{Y}$ are generated under Settings 2.12.4, and we vary $\psi$ over the set $\{0,0.05,0.1,0.15,0.2\}$. For the error terms, we generate $\bm{\epsilon}_{\mathbf{y}}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{y}}})$, and $\bm{\epsilon}_{\mathbf{x}}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{x}}})$.
Setting 2.1.
$\mathbf{X}=\mathbf{Z}\mathbf{B}_{1}+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=\mathbf{Z}\mathbf{B}_{2}+\psi\mathbf{X}\mathbf{B}_{3}+\bm{\epsilon}_{\mathbf{y}}$.
Setting 2.2.
$\mathbf{X}=\mathbf{Z}\mathbf{B}_{1}+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=\sin(\mathbf{Z}\mathbf{B}_{2})+\psi\mathbf{X}\mathbf{B}_{3}+\bm{\epsilon}_{\mathbf{y}}$.
Setting 2.3.
$\mathbf{X}=(\mathbf{Z}\mathbf{B}_{1})^{2}+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=\mathbf{Z}\mathbf{B}_{2}+\psi\mathbf{X}\mathbf{B}_{3}+\bm{\epsilon}_{\mathbf{y}}$.
Setting 2.4.
$\mathbf{X}=\mathbf{Z}\mathbf{B}_{1}+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=\mathbf{Z}\mathbf{B}_{2}+|\psi\mathbf{X}\mathbf{B}_{3}|+\bm{\epsilon}_{\mathbf{y}}$.
The results for Model 2 are shown in Figure 2, with the $x$-axis corresponding to $\psi$. Under $H_{0}$, all six methods effectively control the type-I error. Under $H_{1}$, however, CCIT and CLZ exhibit reduced statistical power. FCIT shows limited statistical power in the linear Setting 2.1 but achieve modest statistical power in the nonlinear Settings 2.22.4, though still under-performing FlowCIT, KCI, and CDC. Our proposed FlowCIT consistently attains the highest statistical power across all settings, with a more pronounced performance gap relative to KCI compared to Model 1. We further investigate high-dimensional scenarios by increasing $(d_{\mathbf{x}},d_{\mathbf{y}},d_{\mathbf{z}})$ in the next section.
Figure 2: Empirical type-I errors and statistical powers of Model 2 with sample size $n=500$. The dashed gray line below refers to significance level of $0.05$. The $x$-axis refers to the value of $\psi$.
4.1.2 Moderately high-dimensional $\mathbf{Z}$
Since CLZ and KCI are kernel-based methods that struggle in high-dimensional settings due to computational burdens, we limit our comparison in Models 34 to CDC, FCIT, and CCIT.
In high-dimensional $\mathbf{Z}$ examples (Models 34), we first consider a low-dimensional pair $(\mathbf{X},\mathbf{Y})$ in Model 3, and then increase the dimensions of $(\mathbf{X},\mathbf{Y})$ in Model 4. In these models, our test demonstrates significantly greater statistical power compared to other methods. While FCIT and CCIT are effective in certain settings, they are ineffective in others.
Model 3 (Low-dimensional $\mathbf{X}$ and $\mathbf{Y}$, and moderately high-dimensional $\mathbf{Z}$).
Let $(d_{\mathbf{x}},d_{\mathbf{y}},d_{\mathbf{z}})=(5,5,50)$ and the sample size $n=1000$. Each element of $\bm{\epsilon}_{\mathbf{x}}$, $\bm{\epsilon}_{\mathbf{y}}$, and $\mathbf{Z}$ is generated from $\text{Normal}(0,1)$. Matrices $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, $\mathbf{B}_{3}$, and vectors $\mathbf{X}$ and $\mathbf{Y}$ are generated under Settings 3.13.4, and the parameter $\psi$ is varied over the set $\{0,0.1,0.2,0.3,0.4\}$.
Setting 3.1 (Row-sparse $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$, and dense $\mathbf{B}_{3}$).
Each element of $\mathbf{B}_{3}$, as well as the first $3\times d_{\mathbf{x}}$ sub-matrix of $\mathbf{B}_{1}$, and the first $3\times d_{\mathbf{y}}$ sub-matrix of $\mathbf{B}_{2}$, is generated from $\text{Normal}(0,1)$. All other elements of $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$ are set to zero. We then generate:
$\mathbf{X}=\mathbf{Z}\mathbf{B}_{1}+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=\mathbf{Z}\mathbf{B}_{2}+\psi\mathbf{X}\mathbf{B}_{3}+\bm{\epsilon}_{\mathbf{y}}.$
Setting 3.2 (Sparse $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$).
Each element of the first $3\times 1$ sub-matrices of $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$ is generated from $\text{Normal}(0,1)$. All other elements of $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{2}$ are set to zero. We then generate:
$\mathbf{X}=\mathbf{Z}\mathbf{B}_{1}+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=(\mathbf{Z}\mathbf{B}_{2})^{2}+\psi\mathbf{X}\mathbf{B}_{3}+\bm{\epsilon}_{\mathbf{y}}.$
Setting 3.3 (Row-sparse $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$, and dense $\mathbf{B}_{3}$).
Each element of $\mathbf{B}_{3}$, as well as the first $2\times d_{\mathbf{x}}$ sub-matrix of $\mathbf{B}_{1}$, and the first $2\times d_{\mathbf{y}}$ sub-matrix of $\mathbf{B}_{2}$, is generated from $\text{Normal}(0,1)$. All other elements of $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$ are set to zero. We then generate:
$\mathbf{X}=\mathbf{Z}\mathbf{B}_{1}+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=\mathbf{Z}\mathbf{B}_{2}+\psi\mathbf{X}\mathbf{B}_{3}+\bm{\epsilon}_{\mathbf{y}}.$
Setting 3.4 (Dense $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$).
Each element of $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$ is generated from $\text{Normal}(0,1)$. We then generate:
$\mathbf{X}=\sin(\mathbf{Z}\mathbf{B}_{1})+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=\mathbf{Z}\mathbf{B}_{2}+|\psi\mathbf{X}\mathbf{B}_{3}|+\bm{\epsilon}_{\mathbf{y}}.$
The results for Model 3 are shown in Figure 3. We find that under sparse settings (Settings 3.13.3), our method demonstrates robust performance, and both FCIT and CCIT also achieve statistical power under the alternative hypothesis. In contrast, in the dense setting (Setting 3.4), our method retains its effectiveness, whereas FCIT and CCIT exhibit no statistical power under the alternative hypothesis.
Figure 3: Empirical type-I errors and statistical powers of Model 3 with sample size $n=1000$. The dashed gray line below refers to significance level of $0.05$. The $x$-axis refers to the value of $\psi$.
Model 4 (Moderately high-dimensional $\mathbf{X}$, $\mathbf{Y}$, and $\mathbf{Z}$).
Let $(d_{\mathbf{x}},d_{\mathbf{y}},d_{\mathbf{z}})=(50,50,100)$ and the sample size $n=1000$. Each element of $\bm{\epsilon}_{\mathbf{x}}$, $\bm{\epsilon}_{\mathbf{y}}$, and $\mathbf{Z}$ is generated from $\text{Normal}(0,1)$. Matrices $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, $\mathbf{B}_{3}$, and vectors $\mathbf{X}$ and $\mathbf{Y}$ are generated under Settings 4.14.4, and the parameter $\psi$ is varied over the set $\{0,0.2,0.4,0.6,0.8\}$.
Setting 4.1 (Row-sparse $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$, and dense $\mathbf{B}_{3}$).
Each element of $\mathbf{B}_{3}$ and the first two rows of $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$ is generated from $\text{Normal}(0,1)$. All other elements of $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$ are set to zero. The sparsity percentages of $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$ are $4\%$. We then generate:
$\mathbf{X}=\mathbf{Z}\mathbf{B}_{1}+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=\mathbf{Z}\mathbf{B}_{2}+\psi\mathbf{X}\mathbf{B}_{3}+\bm{\epsilon}_{\mathbf{y}}.$
Setting 4.2 (Sparse $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$).
Each element of the first $3\times 3$ sub-matrices of $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$ is generated from $\text{Normal}(0,1)$. All other elements of $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{2}$ are set to zero. The sparsity percentages of $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$ are $0.18\%$, $0.18\%$, and $0.36\%$. We then generate:
$\mathbf{X}=|\mathbf{Z}\mathbf{B}_{1}|+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=\mathbf{Z}\mathbf{B}_{2}+\psi\mathbf{X}\mathbf{B}_{3}+\bm{\epsilon}_{\mathbf{y}}.$
Setting 4.3 (Row-sparse $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$, and dense $\mathbf{B}_{3}$).
Matrices $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$ are generated the same as Setting 4.1. We then generate:
$\mathbf{X}=\mathbf{Z}\mathbf{B}_{1}+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=\cos(\mathbf{Z}\mathbf{B}_{2})+\psi\mathbf{X}\mathbf{B}_{3}+\bm{\epsilon}_{\mathbf{y}}.$
Setting 4.4 (Sparse $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$).
Each element of matrices $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$ is randomly taken 0 with probability $90\%$ and generated from $\text{Normal}(0,1)$ with probability $10\%$. The expectation of the sparsity percentages of these matrices are $10\%$. We then generate:
$\mathbf{X}=\cos(\mathbf{Z}\mathbf{B}_{1})+\bm{\epsilon}_{\mathbf{x}},\quad\mathbf{Y}=\mathbf{Z}\mathbf{B}_{2}+\sin(\psi\mathbf{X}\mathbf{B}_{3})+\bm{\epsilon}_{\mathbf{y}}.$
The results for Model 4 are shown in Figure 4. Our method effectively controls the type-I error and demonstrates significantly greater statistical power compared to other methods. FCIT is partially effective in the scenarios with sparse $\mathbf{B}_{1}$, $\mathbf{B}_{2}$, and $\mathbf{B}_{3}$ when the sparsity percentage is below $1\%$ (Setting 4.2). CCIT is partially workable in scenarios with row-sparse $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$ and dense $\mathbf{B}_{3}$ when the sparsity percentages of $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$ are $4\%$ (Setting 4.1 and Setting 4.3). When the sparsity percentage increases to around $10\%$ (Setting 4.4), no method is effective except for our proposed FlowCIT.
Figure 4: Empirical type-I errors and statistical powers of Model 4 with sample size $n=1000$. The dashed gray line below refers to significance level of $0.05$. The $x$-axis refers to the value of $\psi$.
4.2 Application to Dimension Reduction
Conditional independence testing holds broad applicability across statistical methodologies. We demonstrate its practical utility through real-data applications in dimension reduction.
Dimension reduction is a cornerstone of modern statistics and machine learning. Given paired data $(\mathbf{Y},\mathbf{X})$, the goal is to identify a low-dimensional representation $\mathcal{R}(\mathbf{X})$—linear or nonlinear—that satisfies the conditional independence criterion:
$\mathbf{Y}\perp\!\!\!\perp\mathbf{X}\mid\mathcal{R}(\mathbf{X}).$
The efficacy of this reduction hinges critically on both the dimensionality and function form of $\mathcal{R}(\mathbf{X})$. To rigorously evaluate whether a candidate $\mathcal{R}(\mathbf{X})$ achieves sufficiency, we can leverage our conditional independence test framework.
We evaluate dimension reduction methods using the wine quality dataset from Cortez et al. (2009), available at https://archive.ics.uci.edu/dataset/186/wine+quality. The dataset includes $n=4898$ samples of white wine, with predictors (fixed acidity, volatile acidity, citric acid, residual sugar, chlorides, free sulfur dioxide, total sulfur dioxide, density, pH, sulfates, alcohol) and a response variable (wine quality rating).
Our goal is to test whether a dimension reduction method, represented as $\mathcal{R}(\mathbf{X}),$ satisfies the conditional independence $\mathbf{Y}\perp\!\!\!\perp\mathbf{X}\mid\mathcal{R}(\mathbf{X})$. We evaluate four dimension reduction methods:
  1. 1.
    Principal Component Analysis (PCA, Pearson 1901), an unsupervised linear method;
  2. 2.
    Sliced Inverse Regression (SIR, Li 1991), a supervised linear method;
  3. 3.
    Uniform Manifold Approximation and Projection (UMAP, McInnes et al. 2018), an unsupervised nonlinear method;
  4. 4.
    Deep Dimension Reduction (DDR, Huang et al. 2024b), a supervised nonlinear method.
For each method, we test reduced dimensions $d=1$ and $d=4$. The data are split into training (90%) and testing (10%) subsets: $\mathcal{R}(\mathbf{X})$ is learned on the training set, and conditional independence is tested on the held-out set. As a baseline, we set $\mathcal{R}(\mathbf{X})=\mathbf{X}$, which trivially satisfies $\mathbf{Y}\perp\!\!\!\perp\mathbf{X}\mid\mathbf{X}$.
Table 2: $p$-values of conditional independence test of $\mathbf{Y}\perp\!\!\!\perp\mathbf{X}\mid\mathcal{R}(\mathbf{X})$ after learning $\mathcal{R}(\mathbf{X})$ via specific methods. Bold: $p\leq 0.05$; Underlined: $0.05<p\leq 0.1$.
Methods Supervised Nonlinear FlowCIT KCI CDC CCIT FCIT CLZ
PCA-1 0.00 0.04 0.01 0.09 0.00 0.00
PCA-4 0.00 0.00 0.01 0.68 0.27 0.00
SIR-1 $\checkmark$ 0.05 0.09 0.02 0.19 0.94 0.00
SIR-4 $\checkmark$ 0.05 0.00 0.05 0.27 0.79 0.07
UMAP-1 $\checkmark$ 0.00 0.08 0.01 0.29 0.22 0.00
UMAP-4 $\checkmark$ 0.00 0.00 0.01 0.37 0.01 0.00
DDR-1 $\checkmark$ $\checkmark$ 0.00 0.12 0.13 0.32 0.59 0.99
DDR-4 $\checkmark$ $\checkmark$ 0.18 0.14 0.38 0.53 0.82 0.81
$\mathcal{R}(\mathbf{X})=\mathbf{X}$ 0.18 0.52 0.28 0.68 0.18 1.00
Results are summarized in Table 2. Under the baseline ($\mathcal{R}(\mathbf{X})=\mathbf{X}$), all tests fail to reject $H_{0}$, confirming the validity of the testing framework. For PCA, SIR, and UMAP, four tests (FlowCIT, KCI, CDC, and CLZ) reject $H_{0}$ at $\alpha=0.1$, indicating $\mathbf{Y}\not\perp\!\!\!\perp\mathbf{X}\mid\mathcal{R}(\mathbf{X})$ and suggesting these dimension reduction methods inadequately preserve sufficient information. In contrast, the supervised nonlinear DDR method with $d=4$ satisfies $\mathbf{Y}\perp\!\!\!\perp\mathbf{X}\mid\mathcal{R}(\mathbf{X})$ across all tests. In addition, DDR with $d=1$ is rejected by FlowCIT, which suggests that reducing predictors into one dimension is not sufficient.
5 Conclusion
In this work, we propose a novel framework for conditional independence testing that utilizes transport maps. By constructing two transport maps using conditional continuous normalizing flows, we transform the conditional independence problem into an unconditional independence test on the transformed variables. This transformation allows for the application of standard independence measures while preserving statistical validity. Numerical experiments demonstrate the framework’s superior performance across a variety of scenarios, including univariate, low-dimensional, and high-dimensional settings, and demonstrate its robustness to the choice of dependence measures.
In future work, we plan to explore the use of alternative transport maps for converting conditional independence problems into unconditional ones. We will also consider applying the proposed method to other formulations of conditional independence to further enhance the framework’s applicability and effectiveness.
APPENDIX
In the Appendix, we include additional numerical studies and proofs of the theoretical results in the paper.
Appendix A Additional Details of Numerical Studies
A.1 Robustness and Comparison of Dependence Measures
In Section 2.3, we recommend using distance correlation (DC, Székely et al. 2007) and improved projection correlation (IPC, Zhang and Zhu 2024) to evaluate the transformed variables due to their computational efficiency. For both theoretical analysis and numerical studies of our proposed FlowCIT method, we adopt DC as the independence measure. Moreover, FlowCIT exhibits robustness to the choice of independence measures. In this section, we conduct a comparative performance analysis between DC and IPC.
A.1.1 Improved projection correlation
For independence copies $(\bm{\xi}_{1},\bm{\eta}_{1})$, $(\bm{\xi}_{2},\bm{\eta}_{2})$, and $(\bm{\xi}_{3},\bm{\eta}_{3})$, recall the distance covariance defined in (2.7):
$\textup{dcov}^{2}(\bm{\xi},\bm{\eta})=$ $\mathbb{E}\left(\|\bm{\xi}_{1}-\bm{\xi}_{2}\|\|\bm{\eta}_{1}-\bm{\eta}_{2}\|\right)+\mathbb{E}\|\bm{\xi}_{1}-\bm{\xi}_{2}\|\mathbb{E}\|\bm{\eta}_{1}-\bm{\eta}_{2}\|$
$-2\mathbb{E}\left(\|\bm{\xi}_{1}-\bm{\xi}_{2}\|\|\bm{\eta}_{1}-\bm{\eta}_{3}\|\right).$
Improved projection covariance is defined in a similar manner:
$\textup{ipcov}^{2}(\bm{\xi},\bm{\eta})=$ $\mathbb{E}\left\{A(\bm{\xi}_{1},\bm{\xi}_{2})A(\bm{\eta}_{1},\bm{\eta}_{2})\right\}+\mathbb{E}\{A(\bm{\xi}_{1},\bm{\xi}_{2})\}\mathbb{E}\{A(\bm{\eta}_{1},\bm{\eta}_{2})\}$
$-2\mathbb{E}\left\{A(\bm{\xi}_{1},\bm{\xi}_{2})A(\bm{\eta}_{1},\bm{\eta}_{3})\right\},$
where $A(\mathbf{U}_{1},\mathbf{U}_{2})=\arccos\{(\sigma_{u}^{2}+\mathbf{U}_{1}^{\top}\mathbf{U}_{2})(\sigma_{u}^{2}+\mathbf{U}_{1}^{\top}\mathbf{U}_{1})^{-1/2}(\sigma_{u}^{2}+\mathbf{U}_{2}^{\top}\mathbf{U}_{2})\}$ is the arc-cosine function, and $\sigma_{z}^{2}$ is suggested to choose as the median of $\mathbf{U}^{\top}\mathbf{U}$. The improved projection correlation is defined as
$\textup{ipcorr}^{2}(\bm{\xi},\bm{\eta})=\frac{\textup{ipcov}^{2}(\bm{\xi},\bm{\eta})}{\textup{ipcov}(\bm{\xi},\bm{\xi})\textup{ipcov}(\bm{\eta},\bm{\eta})}.$
A.1.2 Comparison in simulations
We conducted simulation studies aligned with the simulation setups of Models 14 to evaluate the sensitivity of FlowCIT to different dependence measures. Results demonstrate that FlowCIT exhibits low sensitivity to the choice of measure, reinforcing its robustness across diverse dependency structures.
As summarized in Table S.1, the empirical type-I errors and statistical powers of FlowCIT-DC and FlowCIT-IPC are nearly identical across all tested scenarios. This consistency highlights that the performance of FlowCIT remains stable irrespective of whether DC or IPC is employed.
Table S.1: Empirical type-I errors and statistical powers under different projection measures used in our FlowCIT. The results suggest that our method is robust to independence measures.
$\psi$ 0 0.1 0.2 0.3 0.4
Model 1 Setting 1.1 DC 0.035 0.350 0.615 0.755 0.835
IPC 0.055 0.345 0.610 0.760 0.840
Setting 1.2 DC 0.050 0.355 0.600 0.745 0.845
IPC 0.055 0.350 0.630 0.765 0.840
Setting 1.3 DC 0.045 0.495 0.790 0.860 0.890
IPC 0.065 0.480 0.775 0.850 0.870
Setting 1.4 DC 0.030 0.490 0.770 0.855 0.885
IPC 0.050 0.495 0.780 0.850 0.880
$\psi$ 0 0.05 0.1 0.15 0.2
Model 2 Setting 2.1 DC 0.055 0.480 0.895 0.990 1.000
IPC 0.055 0.475 0.895 0.985 1.000
Setting 2.2 DC 0.060 0.405 0.925 0.985 1.000
IPC 0.035 0.400 0.930 0.990 1.000
Setting 2.3 DC 0.050 0.590 0.935 1.000 1.000
IPC 0.045 0.535 0.940 0.995 1.000
Setting 2.4 DC 0.065 0.460 0.935 0.995 1.000
IPC 0.050 0.425 0.945 0.995 1.000
$\psi$ 0 0.1 0.2 0.3 0.4
Model 3 Setting 3.1 DC 0.055 1.000 1.000 1.000 1.000
IPC 0.055 1.000 1.000 1.000 1.000
Setting 3.2 DC 0.065 0.225 0.495 0.670 0.820
IPC 0.065 0.190 0.450 0.650 0.780
Setting 3.3 DC 0.055 1.000 1.000 1.000 1.000
IPC 0.025 1.000 1.000 1.000 1.000
Setting 3.4 DC 0.060 0.150 0.365 0.770 0.960
IPC 0.070 0.110 0.190 0.505 0.800
$\psi$ 0 0.2 0.4 0.6 0.8
Model 4 Setting 4.1 DC 0.060 1.000 1.000 1.000 1.000
IPC 0.055 1.000 1.000 1.000 1.000
Setting 4.2 DC 0.055 0.755 0.970 0.995 1.000
IPC 0.045 0.800 1.000 1.000 1.000
Setting 4.3 DC 0.045 1.000 1.000 1.000 1.000
IPC 0.035 1.000 1.000 1.000 1.000
Setting 4.4 DC 0.050 0.340 1.000 1.000 1.000
IPC 0.025 0.260 1.000 1.000 1.000
A.1.3 Comparison in real-data analysis
As shown in Table S.2, we also conduct a comparison between FlowCIT-DC and FlowCIT-IPC in the real-data analysis, and the resulting $p$-values are also very close.
Table S.2: Comparisons of $p$-values between FlowCIT-DC and FlowCIT-IPC in wine-quality dataset. Bold: $p\leq 0.05$.
Methods Supervised Nonlinear FlowCIT-DC FlowCIT-IPC
PCA-1 0.00 0.00
PCA-4 0.00 0.00
SIR-1 $\checkmark$ 0.05 0.05
SIR-4 $\checkmark$ 0.05 0.05
UMAP-1 $\checkmark$ 0.00 0.00
UMAP-4 $\checkmark$ 0.00 0.00
DDR-1 $\checkmark$ $\checkmark$ 0.00 0.00
DDR-4 $\checkmark$ $\checkmark$ 0.18 0.13
$\mathcal{R}(\mathbf{X})=\mathbf{X}$ 0.18 0.16
A.2 Empirical Type-I Error Control
In the main article, we present our results using plots to effectively illustrate the behavior of each method as $\psi$ varies. Since, under $H_{0}$, the data points for these methods are densely clustered, we provide the empirical type-I errors in tabular form in Table S.3. In addition to the methods shown in Figures 14, we include an additional column for FlowCIT-IPC, as discussed in Section A.1.
As shown in Table S.3, the evaluated methods achieve effective type-I error control across a majority of configurations. Mild type-I error inflation is observed in specific cases, such as CDC under the nonlinear settings of Model 2 (Settings 2.22.3), though these deviations remain marginal and within acceptable thresholds.
Table S.3: Empirical type-I errors of Models 14. KCI and CLZ are not included in the comparisons for high-dimensional Models 34 due to their high computational burden.
FlowCIT
-DC
FlowCIT
-IPC
KCI CDC CCIT FCIT CLZ
Model 1 Setting 1.1 0.035 0.055 0.020 0.050 0.005 0.020 0.065
Setting 1.2 0.050 0.055 0.020 0.050 0.005 0.020 0.065
Setting 1.3 0.045 0.065 0.065 0.040 0.005 0.010 0.070
Setting 1.4 0.030 0.050 0.055 0.045 0.000 0.020 0.080
Model 2 Setting 2.1 0.055 0.055 0.055 0.035 0.000 0.045 0.060
Setting 2.2 0.060 0.035 0.030 0.090 0.000 0.035 0.065
Setting 2.3 0.050 0.045 0.080 0.105 0.000 0.060 0.060
Setting 2.4 0.065 0.050 0.055 0.035 0.000 0.050 0.060
Model 3 Setting 3.1 0.055 0.055 0.000 0.005 0.010
Setting 3.2 0.065 0.065 0.000 0.005 0.025
Setting 3.3 0.055 0.025 0.000 0.000 0.010
Setting 3.4 0.060 0.070 0.000 0.000 0.000
Model 4 Setting 4.1 0.060 0.055 0.000 0.060 0.000
Setting 4.2 0.055 0.045 0.000 0.000 0.000
Setting 4.3 0.045 0.035 0.000 0.015 0.030
Setting 4.4 0.050 0.025 0.000 0.000 0.000
A.3 Implementations
Among the compared methods, FlowCIT and CDC require permutation-based method, both configured with $B=100$ permutations. For CDC and FCIT, all other hyperparameters follow the default configurations specified in the Python package hyppo (Panda et al., 2024). The KCI test is implemented using the default settings from the R package CondIndTests::KCI, while CLZ adheres to the original R code. CCIT is executed with the default parameters in the Python package CCIT.
For the FlowCIT, we utilize neural networks to learn the velocity fields. Specifically, we design a ReLU-activated network with two hidden layers: the first layer contains $p_{1}$ neurons, and the second layer reduces dimensionality to $p_{2}=p_{1}/2$ neurons. For low-dimensional conditional independence scenarios (Models 12), we set $p_{1}=32$. In higher-dimensional settings (Models 34), we reduce model complexity by setting $p_{1}=16$ to balance the overall computation units. Full implementation details are publicly available in the GitHub repository: https://github.com/chenxuan-he/FlowCIT.
Appendix B Proof of Lemma 1
In this section, we present the proof for our central idea: transforming the conditional independence structure into an unconditional structure using transport maps, as presented in Lemma 1.
Proof.
The proof for Lemma 1 is given in two steps. We first prove that $\bm{\xi}\perp\!\!\!\perp\mathbf{Z}$ and $\bm{\eta}\perp\!\!\!\perp\mathbf{Z}$, and then show that $\mathbf{X}\perp\!\!\!\perp\mathbf{Y}\mid\mathbf{Z}\iff\bm{\xi}\perp\!\!\!\perp\bm{\eta}$.
Step 1. Prove that $\bm{\xi}\perp\!\!\!\perp\mathbf{Z}$ and $\bm{\eta}\perp\!\!\!\perp\mathbf{Z}$.
In this part, we show that $\bm{\xi}\perp\!\!\!\perp\mathbf{Z}$, and $\bm{\eta}\perp\!\!\!\perp\mathbf{Z}$ can be proven similarly.
Firstly, $\bm{\xi}$ and $\mathbf{Z}$ are independent if their probability density functions satisfy:
$p_{\bm{\xi},\mathbf{Z}}(\mathbf{u},\mathbf{z})=p_{\bm{\xi}}(\mathbf{u})p_{\mathbf{Z}}(\mathbf{z}).$ (S.1)
Based on (S.1), the conditional density of $(\bm{\xi}\mid\mathbf{Z})$ is then given by
$p_{(\bm{\xi}\mid\mathbf{Z})}(\mathbf{u}\mid\mathbf{z})=\frac{p_{(\bm{\xi},\mathbf{Z})}(\mathbf{u},\mathbf{z})}{p_{\mathbf{Z}}(\mathbf{z})}=p_{\bm{\xi}}(\mathbf{u}).$
Thus, to prove the independence between $\bm{\xi}$ and $\mathbf{Z}$, it suffices to show that the law of $(\bm{\xi}\mid\mathbf{Z}=\mathbf{z})$ coincides with the law of $\bm{\xi}$ for all $\mathbf{z}\in\textup{supp}(\mathbf{Z})$. This is satisfied by:
$\bm{\xi}=f(\mathbf{X},\mathbf{Z})\stackrel{{\scriptstyle d}}{{=}}\big{[}f(\mathbf{X},\mathbf{Z})\mid\mathbf{Z}=\mathbf{z}\big{]}.$
Step 2. Prove that $\mathbf{X}\perp\!\!\!\perp\mathbf{Y}\mid\mathbf{Z}\iff\bm{\xi}\perp\!\!\!\perp\bm{\eta}$.
Recall that $\bm{\xi}=f(\mathbf{X},\mathbf{Z})$ and $\bm{\eta}=g(\mathbf{Y},\mathbf{Z})$. Since $\bm{\xi}\perp\!\!\!\perp\mathbf{Z}$ and $\bm{\eta}\perp\!\!\!\perp\mathbf{Z}$, we consider the following probability density functions:
$p_{(\mathbf{X}\mid\mathbf{Z})}(\mathbf{x}\mid\mathbf{z})=$ $p_{(\bm{\xi}\mid\mathbf{Z})}(f(\mathbf{x},\mathbf{z})\mid\mathbf{z})\left|\frac{\partial f(\mathbf{x},\mathbf{z})}{\partial\mathbf{x}}\right|,$
$p_{(\mathbf{Y}\mid\mathbf{Z})}(\mathbf{y}\mid\mathbf{z})=$ $p_{(\bm{\eta}\mid\mathbf{Z})}(g(\mathbf{y},\mathbf{z})\mid\mathbf{z})\left|\frac{\partial g(\mathbf{y},\mathbf{z})}{\partial\mathbf{y}}\right|,$
$p_{(\mathbf{X},\mathbf{Y}\mid\mathbf{Z})}(\mathbf{x},\mathbf{y}\mid\mathbf{z})=$ $p_{(\bm{\xi},\bm{\eta}\mid\mathbf{Z})}(f(\mathbf{x},\mathbf{z}),g(\mathbf{y},\mathbf{z})\mid\mathbf{z})\left|\frac{\partial f(\mathbf{x},\mathbf{z})}{\partial\mathbf{x}}\cdot\frac{\partial g(\mathbf{y},\mathbf{z})}{\partial\mathbf{y}}\right|.$
Since $\bm{\xi}\perp\!\!\!\perp\mathbf{Z}$ and $\bm{\eta}\perp\!\!\!\perp\mathbf{Z}$, it follows that:
$p_{(\mathbf{X}\mid\mathbf{Z})}(\mathbf{x}\mid\mathbf{z})=$ $p_{\bm{\xi}}(f(\mathbf{x},\mathbf{z}))\left|\frac{\partial f(\mathbf{x},\mathbf{z})}{\partial\mathbf{x}}\right|,$ (S.2)
$p_{(\mathbf{Y}\mid\mathbf{Z})}(\mathbf{y}\mid\mathbf{z})=$ $p_{\bm{\eta}}(g(\mathbf{y},\mathbf{z}))\left|\frac{\partial g(\mathbf{y},\mathbf{z})}{\partial\mathbf{y}}\right|,$ (S.3)
$p_{(\mathbf{X},\mathbf{Y}\mid\mathbf{Z})}(\mathbf{x},\mathbf{y}\mid\mathbf{z})=$ $p_{(\bm{\xi},\bm{\eta})}(f(\mathbf{x},\mathbf{z}),g(\mathbf{y},\mathbf{z}))\left|\frac{\partial f(\mathbf{x},\mathbf{z})}{\partial\mathbf{x}}\cdot\frac{\partial g(\mathbf{y},\mathbf{z})}{\partial\mathbf{y}}\right|.$ (S.4)
Thus, since the transport maps $f(\mathbf{X},\mathbf{Z})$ and $g(\mathbf{Y},\mathbf{Z})$ are invertible with respect to the first argument, the Jacobian matrices above are also well-defined. Combining the equations (S.2)–(S.4), we have:
$p_{(\mathbf{X}\mid\mathbf{Z})}(\mathbf{x}\mid\mathbf{z})p_{(\mathbf{Y}\mid\mathbf{Z})}(\mathbf{y}\mid\mathbf{z})=p_{(\mathbf{X},\mathbf{Y}\mid\mathbf{Z})}(\mathbf{x},\mathbf{y}\mid\mathbf{z})\iff p_{\bm{\xi}}(f)p_{\bm{\eta}}(g)=p_{(\bm{\xi},\bm{\eta})}(f,g),$
which implies:
$\mathbf{X}\perp\!\!\!\perp\mathbf{Y}\mid\mathbf{Z}\iff\bm{\xi}\perp\!\!\!\perp\bm{\eta}.$
Appendix C Proofs for the Test Statistic
In this section, we prove the consistency of our test statistic $T_{n}$ stated in Theorem 5 in Section C.1. We prove that the $p$-value obtained by the permutation-based method is reasonable in Section C.2.
C.1 Proof of Theorem 5
Proof.
In the proof, our goal is to demonstrate that the test statistic $T_{n}$ given in (2.12) converges in probability to the true distance correlation $\textup{dcorr}^{2}(\bm{\xi},\bm{\eta})$, i.e., $T_{n}\stackrel{{\scriptstyle p}}{{\to}}\textup{dcorr}^{2}(\bm{\xi},\bm{\eta})$. To achieve this, it suffices to show that $\textup{dcov}_{n}^{2}(\hat{\bm{\xi}},\hat{\bm{\eta}})\stackrel{{\scriptstyle p}}{{\to}}\textup{dcov}(\bm{\xi},\bm{\eta})$, $\textup{dcov}_{n}^{2}(\hat{\bm{\xi}},\hat{\bm{\xi}})\stackrel{{\scriptstyle p}}{{\to}}\textup{dcov}^{2}(\bm{\xi},\bm{\xi})$, and $\textup{dcov}_{n}^{2}(\hat{\bm{\eta}},\hat{\bm{\eta}})\stackrel{{\scriptstyle p}}{{\to}}\textup{dcov}^{2}(\bm{\eta},\bm{\eta})$. These three convergence results are highly similar, and we specifically demonstrate $\textup{dcov}_{n}^{2}(\hat{\bm{\xi}},\hat{\bm{\eta}})\stackrel{{\scriptstyle p}}{{\to}}\textup{dcov}(\bm{\xi},\bm{\eta})$ in the proof.
We first introduce an intermediate statistic $T_{0,n}$:
$T_{0,n}=S_{0,1}+S_{0,2}-2S_{0,3},$ (S.1)
where
$S_{0,1}=$ $\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|\|\bm{\eta}_{k}-\bm{\eta}_{l}\|,$
$S_{0,2}=$ $\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|,$
$S_{0,3}=$ $\frac{1}{n^{3}}\sum_{k=1}^{n}\sum_{l,m=1}^{n}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|\|\bm{\eta}_{k}-\bm{\eta}_{m}\|.$
For the statistic $T_{0,n}$ defined in (S.1), the following Lemma 7 holds.
Lemma 7 (Theorem 2 of Székely et al. 2007).
It holds almost surely that
$\lim_{n\to\infty}T_{0,n}=\textup{dcov}^{2}(\bm{\xi},\bm{\eta}).$
Remark 1.
Since $\bm{\xi}\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}_{0}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{x}}})$ and $\bm{\eta}\stackrel{{\scriptstyle d}}{{=}}\mathbf{Y}_{0}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{y}}})$, the moment conditions for $\bm{\xi}$ and $\bm{\eta}$ required in Székely et al. (2007) are automatically satisfied. This is also an advantage of our method.
Thus, it suffices to prove that:
$\big{|}\textup{dcov}_{n}^{2}(\hat{\bm{\xi}},\hat{\bm{\eta}})-T_{0,n}\big{|}\stackrel{{\scriptstyle p}}{{\to}}0.$
Now we consider the difference between these two statistics, and it follows that:
$|\textup{dcov}_{n}^{2}(\hat{\bm{\xi}},\hat{\bm{\eta}})-T_{0,n}|\leq\underbrace{|S_{0,1}-S_{1}|}_{=:T_{1}}+\underbrace{|S_{0,2}-S_{2}|}_{=:T_{2}}+2\underbrace{|S_{0,3}-S_{3}|}_{=:T_{3}}.$
For the term $T_{1}=|S_{0,1}-S_{1}|$, we rescale it as follows:
$|S_{0,1}-S_{1}|$
$=$ $\bigg{|}\frac{1}{n^{2}}\sum_{k,l=1}^{n}\left(\|\bm{\xi}_{k}-\bm{\xi}_{l}\|\|\bm{\eta}_{k}-\bm{\eta}_{l}\|-\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{l}\|\right)\bigg{|}$
$\stackrel{{\scriptstyle(i)}}{{\leq}}$ $\bigg{|}\frac{1}{n^{2}}\sum_{k,l=1}^{n}\left(\|\bm{\xi}_{k}-\bm{\xi}_{l}\|\|\bm{\eta}_{k}-\bm{\eta}_{l}\|-\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\|\bm{\eta}_{k}-\bm{\eta}_{l}\|\right)\bigg{|}$
$+\bigg{|}\frac{1}{n^{2}}\sum_{k,l=1}^{n}\left(\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\|\bm{\eta}_{k}-\bm{\eta}_{l}\|-\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{l}\|\right)\bigg{|}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}$ $\left(\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|^{2}\right)^{1/2}\left(\frac{1}{n^{2}}\sum_{k,l=1}^{n}\Big{|}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|-\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\Big{|}^{2}\right)^{1/2}$
$+\left(\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|^{2}\right)^{1/2}\left(\frac{1}{n^{2}}\sum_{k,l=1}^{n}\Big{|}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|-\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{l}\|\Big{|}^{2}\right)^{1/2},$ (S.2)
where (i) follows from triangular inequality and (ii) is obtained by Cauchy-Schwarz inequality.
Similarly, for the terms $T_{2}$ and $T_{3}$, we have:
$|S_{0,2}-S_{2}|$
$=$ $\left|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|-\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{l}\|\right|$
$\leq$ $\left|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|-\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|\right|$
$+\left|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|-\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{l}\|\right|$
$\leq$ $\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\Big{|}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|-\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\Big{|}$
$+\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\frac{1}{n^{2}}\sum_{k,l=1}^{n}\Big{|}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|-\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{l}\|\Big{|},$ (S.3)
$|S_{0,3}-S_{3}|$
$=$ $\left|\frac{1}{n^{3}}\sum_{k=1}^{n}\sum_{l,m=1}^{n}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|\|\bm{\eta}_{k}-\bm{\eta}_{m}\|-\frac{1}{n^{3}}\sum_{k=1}^{n}\sum_{l,m=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{m}\|\right|$
$\leq$ $\frac{1}{n^{3}}\sum_{k,l,m=1}^{n}\|\bm{\eta}_{k}-\bm{\eta}_{m}\|\Big{|}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|-\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\Big{|}$
$+\frac{1}{n^{3}}\sum_{k,l,m=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{m}\|\Big{|}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|-\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{l}\|\Big{|}$
$\leq$ $\left(\frac{1}{n^{2}}\sum_{k,m=1}^{n}\|\bm{\eta}_{k}-\bm{\eta}_{m}\|^{2}\right)^{1/2}\left(\frac{1}{n^{2}}\sum_{k,l=1}^{n}\Big{|}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|-\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\Big{|}^{2}\right)^{1/2}$
$+\left(\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|^{2}\right)^{1/2}\left(\frac{1}{n^{2}}\sum_{k,m=1}^{n}\Big{|}\|\bm{\eta}_{k}-\bm{\eta}_{m}\|-\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{m}\|\Big{|}^{2}\right)^{1/2}.$ (S.4)
Now we consider the quantities given in (S.2), (S.3), and (S.4). Since $\bm{\xi}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{x}}})$ and $\bm{\eta}\sim\text{Normal}(\bm{0},\mathbf{I}_{d_{\mathbf{y}}})$, by the law of large numbers, we have:
$\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|^{2}$ $\stackrel{{\scriptstyle p}}{{\to}}2d_{\mathbf{x}},$
$\frac{1}{n^{2}}\sum_{k,l=1}^{n}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|^{2}$ $\stackrel{{\scriptstyle p}}{{\to}}2d_{\mathbf{y}}.$
Thus, it suffices to prove that:
$T_{4}:=\frac{1}{n^{2}}\sum_{k,l=1}^{n}\Big{(}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|-\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{l}\|\Big{)}^{2}$ $\stackrel{{\scriptstyle p}}{{\to}}0,$ (S.5)
$T_{5}:=\frac{1}{n^{2}}\sum_{k,l=1}^{n}\Big{(}\|\bm{\xi}_{k}-\bm{\xi}_{l}\|-\|\hat{\bm{\xi}}_{k}-\hat{\bm{\xi}}_{l}\|\Big{)}^{2}$ $\stackrel{{\scriptstyle p}}{{\to}}0,.$ (S.6)
We prove (S.5), and (S.6) can be done similarly. For the term $T_{4}$, by the triangular inequality, it follows that:
$\frac{1}{n^{2}}\sum_{k,l=1}^{n}\Big{(}\|\bm{\eta}_{k}-\bm{\eta}_{l}\|-\|\hat{\bm{\eta}}_{k}-\hat{\bm{\eta}}_{l}\|\Big{)}^{2}$
$\leq$ $\frac{1}{n^{2}}\sum_{k,l=1}^{n}\Big{(}\|\bm{\eta}_{k}-\bm{\eta}_{l}-\hat{\bm{\eta}}_{k}+\hat{\bm{\eta}}_{l}\|\Big{)}^{2}$
$\leq$ $\frac{1}{n^{2}}\sum_{k,l=1}^{n}\Big{(}\|\bm{\eta}_{k}-\hat{\bm{\eta}}_{k}\|+\|\bm{\eta}_{l}-\hat{\bm{\eta}}_{l}\|\Big{)}^{2}$
$\leq$ $\frac{2}{n}\sum_{k}^{n}\|\bm{\eta}_{k}-\hat{\bm{\eta}}_{k}\|^{2}+\frac{2}{n}\sum_{l}^{n}\|\bm{\eta}_{l}-\hat{\bm{\eta}}_{l}\|^{2}.$
By Lemma 4, we have $\mathbb{E}\|\bm{\eta}-\hat{\bm{\eta}}\|^{2}\stackrel{{\scriptstyle p}}{{\to}}0$. In addition, the estimation error converges to zero under standard regularity conditions in empirical process theory:
$\mathbb{E}_{(\mathbf{X},\mathbf{Z})}\|\hat{\bm{\eta}}-\bm{\eta}\|^{2}-\frac{1}{n}\sum_{i=1}^{n}\big{\|}\hat{\bm{\eta}}_{i}-\bm{\eta}_{i}\big{\|}^{2}\stackrel{{\scriptstyle p}}{{\to}}0\quad\text{ as }n\to\infty.$
C.2 Proof of Theorem 6
Proof.
Recall the original dataset $\mathcal{D}_{n}^{\prime}=\{(\bm{\xi}_{i},\bm{\eta}_{i})\}_{i=1}^{n}$, its estimated version $\mathcal{D}_{n}^{\prime\prime}=\{(\hat{\bm{\xi}}_{i},\hat{\bm{\eta}}_{i})\}_{i=1}^{n}$, the permuted dataset $\mathcal{D}_{n,b}^{\prime}=\{(\bm{\xi}_{i},\bm{\eta}_{\pi_{b}(i)})\}_{i=1}^{n}$, and its estimated version $\mathcal{D}_{n,b}^{\prime\prime}=\{(\hat{\bm{\xi}}_{i},\hat{\bm{\eta}}_{\pi_{b}(i)})\}_{i=1}^{n}$.
Let the permutation $\pi_{b}$ of $\{1,2,\ldots,n\}$ be uniformly random and independent of $\mathcal{D}_{n}^{\prime}$. For the permuted dataset $\mathcal{D}_{n,b}^{\prime}$, the random permutation ensures $(\bm{\eta}_{\pi_{b}(i)}\mid\bm{\xi}_{i})\stackrel{{\scriptstyle d}}{{=}}\bm{\eta}_{\pi_{b}(i)}$, which implies $\bm{\xi}_{i}\perp\!\!\!\perp\bm{\eta}_{\pi_{b}(i)}$ for all $i=1,\ldots,n$. Similar Theorem 5, we obtain
$T_{n,b}\stackrel{{\scriptstyle p}}{{\to}}0.$
Under $H_{0}$, for any fixed threshold $t\in\mathbb{R}^{+}$, this yields
$\mbox{Pr}(T_{n,b}\leq t)-\mbox{Pr}(T_{n}\leq t)\stackrel{{\scriptstyle p}}{{\to}}0.$
Under $H_{1}$, since $T_{n}\stackrel{{\scriptstyle p}}{{\to}}\zeta=\textup{dcorr}^{2}(\bm{\xi},\bm{\eta})>0$, we have
$\mbox{Pr}(T_{n,b}\geq T_{n})=\mbox{Pr}(T_{n,b}\geq\zeta)+o_{p}(1)\stackrel{{\scriptstyle p}}{{\to}}0,$
completing the proof. ∎
Appendix D Proofs and Discussions for the conditional CNFs
D.1 A Discussion of Lemma 3
The convergence result stated in Lemma 3 is a relaxed version of Theorem 5.12 from Gao et al. (2024b). The Lipschitz regularity conditions imposed at Lemma 3 have been proven by Gao et al. (2024b) under several regularity conditions. Based upon the regularity conditions, they proved the convergence:
$\mathbb{E}_{\mathcal{D}_{n}}\mathbb{E}_{(t,\mathbf{X},\mathbf{Z})}\left\|\hat{\mathbf{v}}_{\mathbf{x}}(t,\mathbf{X},\mathbf{Z})-\mathbf{v}_{\mathbf{x}}(t,\mathbf{X},\mathbf{Z})\right\|^{2}\lesssim r_{n},$ (S.1)
with $r_{n}\to 0$ as $n\to\infty$. From this, we can infer that:
$\mathbb{E}_{(t,\mathbf{X},\mathbf{Z})}\left\|\hat{\mathbf{v}}_{\mathbf{x}}(t,\mathbf{X},\mathbf{Z})-\mathbf{v}_{\mathbf{x}}(t,\mathbf{X},\mathbf{Z})\right\|^{2}\stackrel{{\scriptstyle p}}{{\to}}0.$
This inference can be proven by contradiction. Suppose there exists a non-zero probability region where $\mathbb{E}_{(t,\mathbf{X},\mathbf{Z})}\left\|\hat{\mathbf{v}}_{\mathbf{x}}(t,\mathbf{X},\mathbf{Z})-\mathbf{v}_{\mathbf{x}}(t,\mathbf{X},\mathbf{Z})\right\|^{2}$ is bounded away from zero as $n\to\infty$. In that case, after taking expectation over $\mathbb{E}_{\mathcal{D}_{n}}$, $r_{n}$ cannot converge to zero, which contradicts (S.1).
D.2 Proof of Lemma 4
Proof.
Based on Lemma 3, Lemma 4 can be proven by the following inequality:
$\mathbb{E}_{(\mathbf{X},\mathbf{Z})}\|\hat{\bm{\xi}}-\bm{\xi}\|^{2}=\mathbb{E}_{(\mathbf{X},\mathbf{Z})}\|\int_{0}^{1}(\hat{\mathbf{v}}_{\mathbf{x}}-\mathbf{v}_{\mathbf{x}})\textup{d}t\|^{2}$
$\leq$ $\mathbb{E}_{(\mathbf{X},\mathbf{Z})}\int_{0}^{1}\|(\hat{\mathbf{v}}_{\mathbf{x}}-\mathbf{v}_{\mathbf{x}})\|^{2}\textup{d}t=\mathbb{E}_{(t,\mathbf{X},\mathbf{Z})}\|\hat{\mathbf{v}}_{\mathbf{x}}-\mathbf{v}_{\mathbf{x}}\|^{2}\stackrel{{\scriptstyle p}}{{\to}}0,$
where the inequality follows from Jensen’s inequality, since the squared norm is convex. ∎
REFERENCE
  • Albergo et al. (2023) Albergo, M.S., Boffi, N.M., and Vanden-Eijnden, E. (2023). “Stochastic interpolants: A unifying framework for flows and diffusions.” arXiv preprint arXiv:2303.08797.
  • Albergo and Vanden-Eijnden (2022) Albergo, M.S. and Vanden-Eijnden, E. (2022). “Building normalizing flows with stochastic interpolants.” In “The Eleventh International Conference on Learning Representations,” .
  • Bellot and van der Schaar (2019) Bellot, A. and van der Schaar, M. (2019). “Conditional independence testing using generative adversarial networks.” In “Advances in Neural Information Processing Systems,” volume 32. Curran Associates, Inc.
  • Cai and Li (2025) Cai, C. and Li, G. (2025). “Minimax optimality of the probability flow ODE for diffusion models.” arXiv preprint arXiv:2503.09583.
  • Cai et al. (2022) Cai, Z., Li, R., and Zhang, Y. (2022). “A distribution free conditional independence test with applications to causal discovery.” Journal of Machine Learning Research, 23(85), 1–41.
  • Chalupka et al. (2018) Chalupka, K., Perona, P., and Eberhardt, F. (2018). “Fast conditional independence test for vector variables with large sample sizes.” arXiv preprint arXiv:1804.02747.
  • Constantinou and Dawid (2017) Constantinou, P. and Dawid, A.P. (2017). “Extended conditional independence and applications in causal inference.” The Annals of Statistics, 45(6), 2618–2653.
  • Cortez et al. (2009) Cortez, P., Cerdeira, A., Almeida, F., Matos, T., and Reis, J. (2009). “Modeling wine preferences by data mining from physicochemical properties.” Decision Support Systems, 47(4), 547–553.
  • Dawid (1979) Dawid, A.P. (1979). “Conditional independence in statistical theory.” Journal of the Royal Statistical Society: Series B (Methodological), 41(1), 1–15.
  • Duong and Nguyen (2024) Duong, B. and Nguyen, T. (2024). “Normalizing flows for conditional independence testing.” Knowledge and Information Systems, 66(1), 357–380.
  • Fan et al. (2020) Fan, J., Feng, Y., and Xia, L. (2020). “A projection-based conditional dependence measure with applications to high-dimensional undirected graphical models.” Journal of Econometrics, 218(1), 119–139.
  • Fukumizu et al. (2025) Fukumizu, K., Suzuki, T., Isobe, N., Oko, K., and Koyama, M. (2025). “Flow matching achieves almost minimax optimal convergence.” In “The Thirteenth International Conference on Learning Representations,” .
  • Gao et al. (2024a) Gao, Y., Huang, J., and Jiao, Y. (2024a). “Gaussian interpolation flows.” Journal of Machine Learning Research, 25(253), 1–52.
  • Gao et al. (2024b) Gao, Y., Huang, J., Jiao, Y., and Zheng, S. (2024b). “Convergence of continuous normalizing flows for learning probability distributions.” arXiv preprint arXiv:2404.00551.
  • Geiger and Pearl (1993) Geiger, D. and Pearl, J. (1993). “Logical and algorithmic properties of conditional independence and graphical models.” The Annals of Statistics, 21(4), 2001–2021.
  • Ho et al. (2020) Ho, J., Jain, A., and Abbeel, P. (2020). “Denoising diffusion probabilistic models.” In “Advances in Neural Information Processing Systems,” volume 33, pages 6840–6851. Curran Associates, Inc.
  • Huang et al. (2024a) Huang, D., Huang, J., Li, T., and Shen, G. (2024a). “Conditional stochastic interpolation for generative learning.” arXiv preprint arXiv:2312.05579.
  • Huang et al. (2024b) Huang, J., Jiao, Y., Liao, X., Liu, J., and Yu, Z. (2024b). “Deep dimension reduction for supervised representation learning.” IEEE Transactions on Information Theory, 70(5), 3583–3598.
  • Huang et al. (2022) Huang, Z., Deb, N., and Sen, B. (2022). “Kernel partial correlation coefficient — a measure of conditional dependence.” Journal of Machine Learning Research, 23(216), 1–58.
  • Koldanov et al. (2017) Koldanov, P., Koldanov, A., Kalyagin, V., and Pardalos, P. (2017). “Uniformly most powerful unbiased test for conditional independence in Gaussian graphical model.” Statistics & Probability Letters, 122, 90–95.
  • Li (1991) Li, K.C. (1991). “Sliced inverse regression for dimension reduction.” Journal of the American Statistical Association, 86(414), 316–327.
  • Lipman et al. (2022) Lipman, Y., Chen, R.T.Q., Ben-Hamu, H., Nickel, M., and Le, M. (2022). “Flow matching for generative modeling.” In “The Eleventh International Conference on Learning Representations,” .
  • Liu (2022) Liu, Q. (2022). “Rectified flow: A marginal preserving approach to optimal transport.” arXiv preprint arXiv:2209.14577.
  • Liu et al. (2023) Liu, X., Gong, C., and Liu, Q. (2023). “Flow straight and fast: Learning to generate and transfer data with rectified flow.” In “The Eleventh International Conference on Learning Representations,” .
  • McInnes et al. (2018) McInnes, L., Healy, J., Saul, N., and Großberger, L. (2018). “UMAP: Uniform manifold approximation and projection.” Journal of Open Source Software, 3(29), 861.
  • Panda et al. (2024) Panda, S., Palaniappan, S., Xiong, J., Bridgeford, E.W., Mehta, R., Shen, C., and Vogelstein, J.T. (2024). “Hyppo: A multivariate hypothesis testing python package.” arXiv preprint arXiv:1907.02088.
  • Pearl (2009) Pearl, J. (2009). Causality: Models, Reasoning and Inference. Cambridge, U.K. ; New York: Cambridge University Press, 2nd edition edition.
  • Pearson (1901) Pearson, K. (1901). “On lines and planes of closest fit to systems of points in space.” Philosophical Magazine, 2(11), 559–572.
  • Rosenblatt (1952) Rosenblatt, M. (1952). “Remarks on a multivariate transformation.” The Annals of Mathematical Statistics, 23(3), 470–472.
  • Sen et al. (2017) Sen, R., Suresh, A.T., Shanmugam, K., Dimakis, A.G., and Shakkottai, S. (2017). “Model-powered conditional independence test.” In “Advances in Neural Information Processing Systems,” volume 30. Curran Associates, Inc.
  • Shi et al. (2021) Shi, C., Xu, T., Bergsma, W., and Li, L. (2021). “Double generative adversarial networks for conditional independence testing.” Journal of Machine Learning Research, 22(285), 1–32.
  • Song et al. (2020) Song, J., Meng, C., and Ermon, S. (2020). “Denoising diffusion implicit models.” In “International Conference on Learning Representations,” .
  • Song (2009) Song, K. (2009). “Testing conditional independence via Rosenblatt transforms.” The Annals of Statistics, 37(6B), 4011–4045.
  • Spirtes (2010) Spirtes, P. (2010). “Introduction to causal inference.” Journal of Machine Learning Research, 11(54), 1643–1662.
  • Spirtes et al. (1993) Spirtes, P., Glymour, C., and Scheines, R. (1993). Causation, Prediction, and Search, volume 81 of Lecture Notes in Statistics. New York, NY: Springer.
  • Székely et al. (2007) Székely, G.J., Rizzo, M.L., and Bakirov, N.K. (2007). “Measuring and testing dependence by correlation of distances.” The Annals of Statistics, 35(6), 2769–2794.
  • Wang et al. (2015) Wang, X., Pan, W., Hu, W., Tian, Y., and Zhang, H. (2015). “Conditional distance correlation.” Journal of the American Statistical Association, 110(512), 1726–1734.
  • Yang et al. (2024) Yang, Y., Li, S., Zhang, Y., Sun, Z., Shu, H., Chen, Z., and Zhang, R. (2024). “Conditional diffusion models based conditional independence testing.” arXiv preprint arXiv:2412.11744.
  • Zhang et al. (2011) Zhang, K., Peters, J., Janzing, D., and Schölkopf, B. (2011). “Kernel-based conditional independence test and application in causal discovery.” In “Proceedings of the Twenty-Seventh Conference on Uncertainty in Artificial Intelligence,” UAI’11, pages 804–813. Arlington, Virginia, USA: AUAI Press.
  • Zhang and Zhu (2024) Zhang, Y. and Zhu, L. (2024). “Projective independence tests in high dimensions: The curses and the cures.” Biometrika, 111(3), 1013–1027.
  • Zhu et al. (2017) Zhu, L., Xu, K., Li, R., and Zhong, W. (2017). “Projection correlation between two random vectors.” Biometrika, 104(4), 829–843.